### v1

# GraphGhost: Tracing Structures Behind Large Language Models

Abstract
Large Language Models (LLMs) demonstrate remarkable reasoning capabilities, yet the structural mechanisms underlying these abilities remain under explored. In this work, we introduce GraphGhost, a unified framework that represents neuron activations and their signal propagation as graphs, explaining how LLMs capture structural semantics from sequential inputs and generate outputs through structurally consistent mechanisms. This graph-based perspective enables us to employ graph algorithms such as PageRank to characterize the properties of LLMs, revealing both shared and model-specific reasoning behaviors across diverse datasets. We further identify the activated neurons within GraphGhost and evaluate them through structural interventions, showing that edits to key neuron nodes can trigger reasoning collapse, altering both logical flow and semantic understanding. Together, these contributions position GraphGhost as a powerful tool for analyzing, intervening in, and ultimately understanding the structural foundations of reasoning in LLMs.
GraphGhost: Tracing Structures Behind Large Language Models
Xinnan Dai1, Kai Guo1thanks: Correspondence to guokai1@msu.edu.    The code and data are available at https://github.com/DDigimon/GraphGhost, Chung-Hsiang Lo2, Shenglai Zeng1, Jiayuan Ding4, Dongsheng Luo3,Subhabrata Mukherjee4, Jiliang Tang1 1Michigan State University  2 Northeastern University 3 Florida International University  4 Hippocratic AI {daixinna, guokai1}@msu.edu
1 Introduction
Large Language Models (LLMs) have demonstrated promising capabilities in capturing structural information from text Sui et al. (2024); Wang et al. (2025, 2023). Notable examples include identifying relationships between entities in context Zhang and Soh (2024), reasoning and planning through structured thinking processes Yao et al. (2023a); Besta et al. (2024); Wu et al. (2024), and recognizing logical connections across paragraphs Pan et al. (2023). Trained on large-scale corpora, recent studies suggest that LLMs can transfer reasoning processes by extracting semantics and generalizing learned reasoning strategies Muennighoff et al. (2025). This raises the question of whether LLMs internally construct reasoning structures and leverage them across diverse tasks.
Figure 1: An illustration of GraphGhost based on Qwen3-0.6B. a) GraphGhost illustrates a structural mechanism within LLMs that operates consistently across datasets. L indicates the layer number. Sequential inputs activate subsets of neurons, which merge tokens in the intermediate layers to capture semantic meaning. These activated neurons then propagate signals to the logit layers to generate outputs. b) Muting the neuron node. Changing the structure of GraphGhost leads to corresponding changes in both language expression and logical reasoning. “等于” means “equal to” in Chinese.
To understand the mechanisms of constructing reasoning structures from text, prior studies have often simplified reasoning tasks into toy examples such as graph-based reasoning to examine how LLMs trace reasoning paths within giving structures Sanford et al. (2024). The underlying hypothesis is that uncovering how decoder-only transformers construct such graphs may yield insights that generalize to real-world tasks Saparov et al. (2025). While these toy graph settings provide preliminary insights—highlighting mechanisms such as token merging Dai et al. (2025b) and edge memorization Wang et al. (2024), they fall short when applied to complex, real-world structured data. To bridge this gap, we introduce GraphGhost, a framework that provides a structural perspective on how LLMs across different architectures perform reasoning tasks in diverse domains. GraphGhost captures neurons and their signal propagation inside LLMs, revealing how neurons are activated across samples and datasets, enabling shared semantic understanding and supporting the construction of reasoning processes across contexts.
As illustrated in Figure 1 a), GraphGhost represents a structural mechanism modeled within LLMs, where neuron activations capture semantic information and guide token outputs. In the sentence-reading stage, LLMs merge tokens to derive semantic meaning, thereby activating specific neurons associated with those tokens. These activated neurons then propagate signals to higher layers, ultimately influencing the logit neurons in the output layers. Notably, while certain neurons are activated for the given samples, many remain silent. These silent neurons are reserved for unseen inputs from different domains, enabling shared output behaviors across contexts. For example, the symbol ‘X’ may activate the ‘=’ symbol in mathematical equations, even if the current sample does not require it, leaving those neurons in a silent state until needed.
We define neurons as nodes and the signal propagation among neurons as edges. This representation allows us to analyze LLMs from a graph perspective. For instance, by inspecting high in-degree nodes or applying PageRank, we can identify key nodes that control next-token prediction, as further elaborated in Section 3. Furthermore, we evaluate the effectiveness of these nodes through structural interventions, such as node removal (Figure 1 b)). Removing specific nodes can lead to semantic shifts, such as switching to another language, or logical changes. In addition, we conduct statistical analyses of these structural perturbations to quantify their effects, as presented in Section 4. In conclusion, our key contributions are summarized as:
  1. 1.
    We introduce GraphGhost, a unified framework represented in graphs, explaining how LLMs capture structural semantics from sequential inputs and generate outputs through a structurally consistent mechanism.
  2. 2.
    We provide a graph-based perspective for understanding LLMs and apply graph algorithms on GraphGhost to characterize the properties of LLMs, revealing both unique and shared behaviors across different models and datasets.
  3. 3.
    We identify the activated neurons within LLMs and evaluate them by editing the structure of GraphGhost. Our experiments show that key neuron nodes can trigger reasoning collapse, altering both the logical flow and semantic understanding.
2 Related work and Preliminary
GraphGhost is built on circuit tracing, which is a typical LLM interpretation tool. In this section, we first introduce the background on LLM interpretability and reasoning tasks in Section 2.1 and Section 2.2, respectively. To illustrate how LLMs construct internal structures that align with explicit reasoning texts, we design a toy experiment in Section 2.3 to illustrate the alignment of inner signal propagation with explicit structure descriptions. Building on this foundation, we provide the preliminary definition in Section 2.4.
2.1 Interpretation of LLMs
Interpreting how LLMs perform reasoning is essential for aligning their behavior with human understanding Zhao et al. (2024). Existing approaches fall into two categories, self-explanation Gu et al. (2024) and causal-mechanistic interpretation Bloom et al. (2024). The details are suggested in Appendix A.1
The causal-mechanism approach focuses on explaining how each token is predicted Anthropic (2023); Ameisen et al. (2025). Based on learning the internal reasoning pathways, Lindsey et al. (2025) introduced attribution graphs to capture token flows that contribute to final predictions. Formally, an attribution graph is defined as $G_{A}=(\hat{V},\hat{E},A)$, where $\hat{V}$ denotes activated neurons, $A_{s\to t},;s,t\in\hat{V}$ represents the contribution from node $s$ to $t$ (computed via backward Jacobians), and $\hat{E}$ indicates whether an edge exists. The edge set is controlled by an edge ratio, which filters the top influential connections according to $A$.
2.2 Reasoning Tasks
Reasoning tasks including mathmatical Cobbe et al. (2021), scientific Khot et al. (2020), and logic reasoning Clark et al. (2019) are set to evaluate the ability in solving complex problems. The details are in Appendix C.2
Graph reasoning represents a distinct type of reasoning task Dai et al. (2025a). In this setting, models are provided with textual descriptions of graph structures and asked to reason over them. Success on such tasks suggests that LLMs can leverage implicit structural representations to guide reasoning Wang et al. (2023). However, how these implicit structures are internally constructed and maintained within LLMs remains an open question.
2.3 Path tracing in the decoder-only transformer
Graph reasoning tasks, particularly path reasoning, offer a prototypical setting for understanding how structural information in sequential inputs can be mapped to outputs that describe those structures Wang et al. (2023). In this task, the model is provided with a graph description and a query consisting of a start node and an end node, and the objective for LLMs or decoder-only transformers is to predict the path connecting the two nodes Dai et al. (2024); Wang et al. (2024). We adopt this basic setup to investigate how decoder-only transformers perform path reasoning. Following prior work Wang et al. (2024), we use simplified graph descriptions, and a query (“from start point S to end point E”) and must predict the connecting path. as illustrated in Figure 2(a), and train a 5-layer GPT-2 architecture decoder-only transformer on this task. The trained model achieves 98% accuracy on the test set. We then apply the circuit-tracing method, specifically using transcoder Dunefsky et al. (2024), to analyze how the transformer models the path-finding process, as shown in Figure 2(b). Full experimental details are provided in Appendix C.1.
Circuit tracing reveals how decoder-only transformers review the previous context to identify related nodes and predict the next token. From a structural understanding perspective, neighboring nodes are merged in the middle layers. For example, at the third layer, node 4 is merged with node 5, forming the semantic representation of the edge between them. Further, in generating the output tokens, the final token in the sequence receives the message that an edge exists between 4 and 5 at layer 3, leading the model to predict 5 as the next token.
From this toy example, we observe that decoder-only transformers are not only capable of reasoning about paths within given graph structures, but that their underlying mechanism involves activated neurons propagating signals across layers. However, beyond the toy example with decoder-only transformers, does this mechanism extend to LLMs that can solve more complex tasks expressed in natural language? To address this, we introduce GraphGhost, a unified framework that represents these neuron activations and their signal flows as graphs. GraphGhost enables us to trace and mergeattribution graphs defined in Section 2.1 across diverse reasoning tasks, revealing how LLMs encode and reuse structural reasoning patterns. In the following sections, we apply GraphGhost to multiple models and domains to analyze how their internal reasoning structures build up, interact, and generalize.
(a) Graph description in a simplified way. The sequence includes the graph description, query, and answer.
(b) Interpretation of the graph path reasoning task. The bolder lines indicate stronger contributions to neuron activations.
Figure 2: A toy example of path reasoning on a graph. (a) The simplified descriptions to predict the path between 4 and 7; (b) The circuit tracing for the interpretation of why token 5 is selected to be predicted.
2.4 Preliminary definitions
We formalize the core concepts underlying GraphGhost. GraphGhost is constructed from attribution graphs obtained using $\text{CircuitTracer}(\cdot)$, which traces neuron activations contributing to next-token prediction and records their signal flow. Let $X=(t_{1},\dots,t_{n})$ be the input tokens of an LLM with layers $\{f^{(1)},\dots,f^{(L)}\}$. For a target token $t_{n+1}$, the attribution graph is defined as $A_{n+1}=\text{CircuitTracer}(t_{n+1}),$ where each node represents an activated neuron $(t_{i},\ell)$ with activation $h^{(\ell)}(t_{i})=f^{(\ell)}(h^{(\ell-1)}(t_{i}))$, and each directed edge $w_{s\to t}$ encodes the contribution from neuron $s$ to $t$. The set of contributing tokens is $\hat{T}(t_{n+1})=\{\,t_{i}\mid(t_{i},\ell)\in A_{n+1}\,\}$. where $\ell$ denotes the layer index. These local attribution graphs form the building blocks of GraphGhost, which aggregates them across samples into a unified weighted structure.
Following the definition of attribution graphs in Ameisen et al. (2025), a neuron node is represented as a pair $(t_{i},\ell)$, where $t_{i}$ denotes a token and $\ell$ denotes the layer index. Each neuron node has a corresponding hidden activation $h^{(\ell)}(t_{i})$, representing the hidden state of token $t_{i}$ at layer $\ell$, computed as $h^{(\ell)}(t_{i})=f^{(\ell)}\big(h^{(\ell-1)}(t_{i})\big).$ For a source neuron node $s=(t_{i},\ell_{s})$ and a target neuron node $t=(t_{j},\ell_{t})$ with $\ell_{s}<\ell_{t}$, the edge weight in the attribution graph is defined as $w_{s\to t}=A_{s\to t},$ where $A_{s\to t}$ denotes the attribution score from $s$ to $t$.
For example, in Figure 2,to generate token $5$, the contributing tokens are $\{4,5,7,\texttt{"|"},\texttt{"S"},\texttt{"P"}\}$. The corresponding neuron nodes include $h^{(3)}(5)$, $h^{(4)}(7)$, and $h^{(5)}(4)$, among others. The edge weight from $h^{(3)}(5)$ to $h^{(5)}(4)$ is denoted as $w_{h^{(3)}(5)\to h^{(5)}(4)}$.
3 GraphGhost in LLMs
Beyond the toy examples, we apply GraphGhost to real-world LLMs to uncover their internal reasoning structures. GraphGhost traces neuron activations that contribute to next-token generation, represents them as local attribution graphs, and merges these across samples into a global weighted graph. This unified structure reveals stable signal propagation patterns shared across tasks and models. Section 3.1 details the tracing and aggregation process. To evaluate whether these structures correspond to meaningful reasoning behaviors, we introduce the experimental setup in Section 3.2. which specifies the models, datasets, and tracing configurations used for constructing GraphGhost. Section 3.2 and Section 3.4 then present case studies and quantitative analyses based on these settings.
3.1 Tracing the GraphGhost
To construct a unified GraphGhost that captures reasoning structures shared across multiple samples, we aggregate the attribution graphs obtained from each sample. Given a set of samples $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$, where each sample $s_{i}$ contains an answer token $t_{n}^{(i)}$, we initialize a global graph $G=(V,E,W)$ with empty node, edge, and weight sets. For each sample $s_{i}$, we start from its answer token $t_{n}^{(i)}$ and iteratively trace the contributing tokens and their local attribution graph using the $\text{CircuitTracer}(\cdot)$ function. Each traced graph produces a set of neuron nodes $\hat{V}$ and directed edges $\hat{E}$, which are merged into the global GraphGhost. If an edge $e\in\hat{E}$ does not exist in $E$, it is added with an initial weight $W(e)=1$; otherwise, its weight is incremented by one. This cumulative weighting reflects the frequency of edge occurrences across samples, capturing the stability and recurrence of signal propagation pathways. After processing all samples, the resulting GraphGhost $G=(V,E,W)$ represents a multi-sample structural abstraction of the model’s internal reasoning dynamics, which is a weighted directed graph describing neuron-level signal propagation. The full procedure is summarized in Algorithm 1 fromAppendix B.
3.2 Experiments setting
We evaluate a range of LLMs and datasets to investigate whether GraphGhost constitutes a general mechanism across models. Specifically, we consider Qwen3-0.6B (28 layers), Qwen2.5-1.5B (28 layers), Llama3.1-8B-Instruct (32 layers), DeepSeek-distilled-Qwen2.5B-1.5B (DS-Qwen-1.5B) (28 layers), and DeepSeek-distilled-Llama-8B (DS-Llama-8B) (32 layers). For datasets, we examine three domains: (i) mathematical reasoning (GSM8K, MAWPS), (ii) logical reasoning (ProntoQA, BoolQA), and (iii) scientific reasoning (ARC-Easy, QASC). To elicit the models’ reasoning processes, we apply Chain-of-Thought prompting. Following Hanna et al. (2025), we then train transcoders at every layer for each dataset–model pair to probe neuron activations during reasoning, ensuring that the transcoders effectively capture the most strongly activated neurons. All the experiments are run on a single H200 GPU. Full experimental details are provided in Appendix C.2.
3.3 Case study on GraphGhost
Figure 3: Case study of tracing contributing tokens for the answer 14 in Qwen3-0.6B model. We highlight the numerical reasoning flow captured by GraphGhost to visualize how number-related tokens contribute to the final prediction.
We further provide some case studies of GraphGhost on contributing nodes and substructures, respectively. Figure 3 is an illustrative example of tracing contributing tokens within a single sample. The identified token set $T$ includes tokens from both the input question (highlighted in green) and those generated along the reasoning chain (in orange). As shown in the figure, number-related tokens emerge through three distinct mechanisms: (i) step-by-step progression (e.g., 1–2–3–4), (ii) direct copying from the input context (e.g., 30 and 28), and (iii) generation influenced by neighboring tokens, such as the digit “1” in the final answer “14”, which is activated by “=". This tracing process demonstrates how local dependencies between tokens collectively form reasoning paths during sequential generation.
Figure 4 demonstrates a representative subgraph extracted from the GraphGhost constructed on Qwen3-0.6B model, integrating neuron activations from mathematical reasoning (GSM8K and MAWPS), scientific reasoning (ARC), and logical reasoning (ProntoQA). Each neuron node is annotated with its corresponding token and layer ID. For example, the activated period at layer 22 is denoted as $_{22}.$ This figure highlights how neuron activations from diverse domains converge into a unified structural representation, revealing consistent signal propagation patterns underlying different types of reasoning tasks.
Figure 4: A subgraph of GraphGhost in Qwen3-0.6B model. We include 4 datasets from different domains.
In the subgraph, we observe that some nodes are activated consistently across datasets, while others contribute only to specific domains or even individual samples. For example, the period symbol “.” at layer 22 triggers strong logit activations across all four datasets, and the token “So” frequently appears across different reasoning types. Domain-specific reasoning tokens also emerge within the same domain. For instance, the action “Adding” and arithmetic expressions such as “6+” appear in both GSM8K and MAWPS. In contrast, certain nodes are unique to individual samples—for example, $\mathrm{B_{18}}$ activates only in choice-option reasoning within the ARC dataset.
3.4 Analysis on GraphGhost
In Figure 4, we also observe two typical actions associated with the node $_{22}.$ The first is semantic understanding, exemplified by the “6+” operation. The second is logic shifting, where the activation leads to the token “So” in the subsequent prediction. We further analyze these two actions from a statistical perspective.
Semantic understanding
In Section 2.3, we discussed how edge information is merged within the inner layers. Building on this, we analyze the token-merging behavior of neuron nodes by examining their average in-degree across layers in LLMs’ settings. The statistical results, summarized in the Appendix D and illustrated in Figure 5, show the top tokens across different datasets and models. We observe that the tokens with the highest in-degree are domain-sensitive. For example, in the ProntoQA dataset, the model is more sensitive to the period token than to the comma. This suggests that LLMs activate specific neurons in response to domain-specific data to capture semantics. Across models, however, the activated neurons are generally consistent: tokens such as <space>, period, and “the" consistently appear among the top-ranked, and this alignment persists even after finetuning.
Figure 5: The top merging nodes across the models and datasets.
We further analyze the merging behavior across layers by computing the in-degree ratio. Let $d_{i}^{(\ell)}$ denote the in-degree of neuron node $(t_{i},\ell)$, where $\ell$ is the layer index and $i$ indexes the neuron within that layer. The in-degree ratio of layer $\ell$ is then defined as $R_{\ell}=\frac{\sum_{i=1}^{N_{\ell}}d_{i}^{(\ell)}}{\sum_{k=1}^{L}\sum_{i=1}^{N_{k}}d_{i}^{(k)}},$ where $N_{\ell}$ is the number of neurons in layer $\ell$ and $L$ is the total number of layers. As shown in Figure 6, the in-degree ratio distribution is imbalanced across both models and tokens. Although Qwen models and the Llama model perform semantic merging primarily in the bottom and top layers, the exact layers differ between them. Similarly, token merging patterns are not always consistent. Most tokens are merged in shallow layers, whereas high in-degree tokens such as period and “the" tend to merge at higher layers.
Figure 6: The in-degree proportions on various layers with MAWPS dataset
Logic shifting.
The reasoning structure is segmented by special tokens such as So, Step, and Wait Muennighoff et al. (2025); Bogdan et al. (2025). We assume that activated neuron nodes in the middle layers trigger these special token expressions at the logit layers. At the same time, the neuron nodes themselves are activated by input tokens. To capture both effects, we treat logit-layer nodes as signals directed back to the neuron nodes by reversing their edge directions. We then apply the PageRank algorithm to identify the most influential nodes. We summarize the results in Figure 7 and Figure 8, and the full details can be found in Appendix D.
Figure 7: Top contributing tokens and layers in Qwen3-0.6B in various datasets.
Figure 8: Top contributing tokens and layers across models evaluated by MAWPS dataset.
Figure 7 presents the token distribution across datasets, highlighting the distinction between silent and activated tokens in LLMs when processing data from different domains. For example, numerical tokens such as 0 are more frequently activated in mathematical datasets, while general tokens like the period and “the" appear across multiple domains. Furthermore, Figure 8 illustrates that different models exhibit distinct activation preferences. After finetuning with distilled data, DS-Qwen-1.5B and DS-Llama-8B show increased activation of the token "**", suggesting that DeepSeek-distilled data encourages reliance on this token during generation. From a layer-level perspective, activated nodes consistently emerge in higher layers across both models and datasets, though not exclusively in the final layer; in the Qwen series, layers 22 to 27 frequently contain critical activated nodes.
4 GraphGhost affects the reasoning process
GraphGhost paves a new way in which LLM reasoning processes can be analyzed and controlled. To evaluate this, we perform perturbation experiments at two complementary levels. (i) At the input level, we modify the contributing tokens identified in Section 4.1 to examine how altering key linguistic elements affects the model’s reasoning trajectory and output behaviors. (ii) At the neuron level, we intervene on the influential neuron nodes detected via in-degree analysis and PageRank in Section 4.2, assessing how structural disruptions within the reasoning pathways influence the model’s internal logic flow. Together, these two perturbation perspectives reveal how surface-level token changes and internal structural shifts jointly shape the reasoning dynamics captured by GraphGhost.
4.1 Impact of contribution tokens
GraphGhost in LLMs leverages contributing tokens to capture context and perform reasoning. In practice, we assess their importance by masking tokens, replacing each with a blank token. To quantify this effect, we evaluate two metrics: Sparsity and Fidelity. Sparsity measures the proportion of contributing tokens in a given question, while Fidelity evaluates whether the answer generated from only the contributing tokens remains consistent with the original answer. Formally, $\mathrm{Sparsity}=\frac{\#C}{\#T}$ and $\mathrm{Fidelity}_{T}=\mathds{1}(LLM(C),LLM(T))$, where $C$ denotes the set of contributing tokens, $T$ is the full set of tokens, and $\mathds{1}$ is an indicator function returning 1 if the two answers are identical and 0 otherwise. In practice, we use gemini-2.5-flash-lite to identify whether the two predictions are the same, with the prompt as: "Comparing two answers for a question: Answer1 {LLM(T^)LLM($\hat${T})}, Answer2 {LLM(T^)LLM($\hat${T})}. Do they have the same final results?" To compare the effectiveness of GraphGhost across models, we evaluate 100 samples that are solvable by Qwen3-0.6B, Qwen2.5-1.5B, and DS-Qwen-1.5B. The results are presented in Figure 9.
Figure 9: Sparsity Vs. Fidelity in various datasets and models.
Change the semantic Change the language Question: A ) nervous and digestive B ) nervous and circulatory C ) respiratory and digestive D ) respiratory and circulatory Question: Which two body systems would most directly remove extra fluid from a person’s lungs? Original answer: … B) Nervous and circulatory systems: The nervous system is responsible for regulating the body’s response to stress, while the circulatory system is responsible for transporting blood throughout the body. Neither of these systems directly remove extra fluid from the lungs… Mute the neuron node ’the’ at layer 20 Perturbed answer: … B) Nervous and circulatory systems: The nervous system is responsible for controlling the - circulatory system is responsible for transporting blood throughout the分区肺. While the分区肺需要血液来清除多余的水分,这些系统本身并不直接参与这个过程。…   Change the word expression Question: To determine how many trips Jerry will make, we need to calculate the total number of trays he needs to pick up and then divide by the number of trays he can carry at a time. Original answer: … we divide the total number of trays, 16.0 trays ÷ 8.0 trays per trip = 2 trips… Mute the neuron node ’.’ at layer 17 Perturbed answer: … Number of trips = Total trays / 8 = 16 / 8 = 2 trips …
Figure 10: Changing the semantic understanding. The characters in changing the language case is translating the original answer.
As shown in Figure 9, compared with random masking of the given question, the contributing tokens identified by GraphGhost successfully preserve the generated answer. This effect is particularly significant in mathematical reasoning tasks such as MAWPS and GSM8K, where even when half of the question tokens are removed, LLMs are still able to produce the correct prediction. In contrast, the performance gap between randomly masked and selected tokens is smaller in scientific reasoning tasks such as ARC dataset. This is because in ARC the answers often involve selecting the correct option, and in some cases LLMs rely on direct recognition of the options from the given context rather than explicit reasoning.
4.2 Modification on the graph structure
To evaluate whether GraphGhost successfully identifies the most influential tokens in LLMs, we mute specific neuron nodes and observe the resulting changes in the outputs. Specifically, neuron nodes are muted by setting the corresponding positions of their layer–token pairs to zero during the reasoning process. We evaluate the impact through both case studies and statistical analysis.
Case study
First, as discussed in Section 3.4, the impacts include changes in semantic understanding and logic shifting. As shown in Figure 10, the semantic understanding is reflected in changing the language formulation and word expression. When the neuron node “the” is deleted at layer 20, the English semantic pathway is disrupted, activating Chinese neurons that share the same semantic signal. As a result, the output appears in Chinese, directly translated from English. Furthermore, in the word-expression change case, deleting the neuron “.” at layer 17 causes the token “/” to replace “÷.” This suggests that another neuron is responsible for propagating the division signal, expressing it as “/.” These findings indicate that LLMs store semantics redundantly across multiple neuron nodes: when the most relevant neuron is muted, the model relies on alternative neurons that produce similar expressions. In Figure 11, we demonstrate how GraphGhost influences reasoning by altering key steps or driving the generation into loops. In the key-step change case, deleting the neuron node “.” at layer 25 introduces a new reasoning token such as “So", ultimately leading the model to produce the correct answer. In the loop case, deleting the neuron node “the” at layer 27 causes the reasoning to degenerate into a repetition, producing “theShaper Information.” These cases suggest that even a single neuron in the middle layer can fundamentally change the outcome of the reasoning process.
Change the reasoning structure Change the key steps Question: There are 7.0 crayons in the drawer and 6.0 crayons on the desk. Sam placed 4.0 crayons and 8.0 scissors on the desk. How many crayons are now there in total? Let’s think step by step (ground truth: 17) Original answer: … - Sam placed 4.0 crayons on the desk. - Sam also placed 8.0 scissors on the desk. …. - Total crayons on the desk = 7.0 + 4.0 + 8.0 - Total crayons on the desk = 19.0 Mute the neuron node ’.’ at layer 25 Perturbed answer: … - Sam placed 4.0 crayons on the desk. - So, the total number of crayons on the desk is now 6.0 (initial crayons) + 4.0 (Sam’s crayons) = 10.0 crayons. … - So, the total number of crayons = 7.0 (drawer) + 10.0 (desk) = 17.0 crayons.   Failed into loop Question: Alyssa picked 25.0 limes and Mike picked 32.0 limes. Tom picked 12.0 plums. How many limes were picked in all? Let’s think step by step Original answer: Step 1: Identify the number of limes picked by Alyssa and Mike. Alyssa picked 25.0 limes. Mike picked 32.0 limes. … Mute the neuron node ’the’ at layer 27 Perturbed answer: Step 1: Identify theSharper Information In this problem, we are given theSharper Information about theSharper Information about Alyssa, Mike, and Tom’s limes and plums. We need to find theSharper Information about theSharper Information about theSharper …
Figure 11: Changing on the logic shifting
Statistical analysis
[…]
We first evaluate the tokens ranked by the PageRank algorithm. Specifically, we calculate the average InFidelity score when muting the corresponding neuron nodes and identify the maximum-impact neuron nodes within the top-$k$ positions. The results, summarized in Table 1, include the top-1–10 and top-11–20 nodes. Overall, the top-10 nodes show a markedly higher influence on the model’s reasoning process, indicating that the PageRank algorithm effectively identifies influential nodes in the GraphGhost framework. Moreover, larger models demonstrate greater stability than smaller ones—for instance, Llama-8B achieves an InFidelity score of 0.28, whereas Qwen3 reaches 0.58. Interestingly, the <space> token emerges as the most impactful token, exerting a strong influence on overall model performance, particularly in higher layers.
Top1-10 Top11-20
AVG MAX AVG MAX
Token - <space>_27 - 0_23
Qwen3-0.6B 0.5845 0.9701 0.3756 0.6958
Token - <space>_27 - <space>_31
Qwen2.5-1.5B 0.4586 0.9577 0.3411 0.8956
Token - <space>_31 - <space>_30
Llama-8B 0.2808 0.8380 0.2454 0.5028
Table 1: The impacts of top tokens. “MAX” denotes the maximum influential token among the top-K slots
We further analyze the relationship between the in-degree ratio across layers and the final reasoning predictions in the Qwen3-0.6B model, as illustrated in Figure 12. Neuron nodes with higher in-degree values are likely to merge semantic information from multiple sources. To verify this, we compute the language-changing ratio when muting individual neuron nodes across layers, defined as $R=\frac{\#\mathrm{non-English\ answer}}{\#\mathrm{All\ answer}}$.where a non-English answer refers to any generated response that switches from English to another language after perturbation. As shown in Figure 12(a), the peaks of the language-changing ratio align closely with layers exhibiting high in-degree ratios. These layers are primarily distributed in the upper regions of the model, which also correspond to the increase in InFidelity scores observed in Figure 12(b). This pattern suggests that neuron nodes in higher layers are more likely to connect directly to logit neurons, thereby altering the logical flow of reasoning (e.g., causing generation loops) rather than simply substituting equivalent semantic representations.
Figure 12: The impact of tokens on various layers. We mute the period token ’.’ for each layer.
5 Conclusion
In this work, we introduced GraphGhost, a framework that models neuron activations and their propagation as graphs, providing a structural view of LLM reasoning. Using graph algorithms and structural interventions, we show how key neurons shape the logical flow and semantic consistency. These findings establish GraphGhost as an interpretation framework for understanding the structural foundations of reasoning in LLMs.
Limitations
The fidelity of GraphGhost relies heavily on the precision of the underlying attribution tracing method. Because current attribution estimates can be noisy or incomplete, they may produce missing or spurious edges, yielding only partial views of the model’s reasoning process. Advancing the design of circuit tracer would further enhance GraphGhost’s ability to identify key neuron nodes and capture more precise reasoning structures.
Appendix A Related work
A.1 Interpretation
Interpreting how LLMs perform various reasoning tasks is an important step toward aligning them with human understanding Yao et al. (2023b); Zhao et al. (2024). Current work mainly falls into two directions: self-explanation Gu et al. (2024); Koa et al. (2024) and uncovering the underlying causal mechanisms Bloom et al. (2024); Dunefsky et al. (2024). The self-explanation approach, such as Chain-of-Thought prompting Wei et al. (2022); Kojima et al. (2022), requires LLMs to verbalize their reasoning steps, which can help guide them to correct answers. However, there is ongoing debate about whether such reasoning chains are truly faithful Turpin et al. (2023); Barez et al. (2025).
A.2 Reasoning tasks
Reasoning tasks are among the most important benchmarks for evaluating the ability of LLMs to solve complex problems. Specifically, they assess model performance across mathematical, scientific, and logical reasoning. Mathematical reasoning evaluates a model’s capacity to perform structured, multi-step numerical computations expressed in natural language Cobbe et al. (2021); Hendrycks et al. (2021). Scientific reasoning tasks require LLMs to draw on background knowledge and apply it in reasoning processes Khot et al. (2020); Clark et al. (2018). Logical reasoning tasks assess whether models can apply formal rules rather than relying on surface-level correlations Clark et al. (2019); Saparov and He (2022). Recent research further suggests that LLMs are capable of transferring reasoning strategies across diverse tasks, indicating that despite differences in domain, the underlying reasoning logic may be shared.
Appendix B Algorithm
Here we present the algorithm of construct the GraphGhost in Algorithm 1.
Algorithm 1 GraphGhost Construction
1:Input: A set of samples $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$, where each $s_{i}$ contains an answer token $t_{n}^{(i)}$
2:Initialize global GraphGhost $G=(V,E,W)$, $V=\emptyset$, $E=\emptyset$, $W=\emptyset$
3:for each sample $s_{i}\in\mathcal{S}$ do
4:  Initialize token list $T=\{t_{n}^{(i)}\}$
5:  for each token $t\in T$ do
6:   if Token $t$ not in question of $s_{i}$ then
7:     # trace contributing tokens and local attribution graph
8:     $\hat{T},G_{A}^{\prime}\leftarrow\text{CircuitTracer}(t)$
9:     $T\leftarrow T\cup\hat{T}$
10:     # update global GraphGhost with local graph
11:     $\hat{V}=\{v^{1}(t_{1}),\dots,v^{\ell}(t_{n})\},\hat{E}\leftarrow G_{A}^{\prime}$
12:     $V\leftarrow V\cup\hat{V}$
13:     for each edge $e$ in $\hat{E}$ do
14:      if $e\notin E$ then
15:        $E\leftarrow E\cup\{e\}$
16:        $W(e)\leftarrow 1$
17:      else
18:        $W(e)\leftarrow W(e)+1$
19:      end if
20:     end for
21:   end if
22:  end for
23:end for
24:return $G=(V,E,W)$
Appendix C Appendix: Experiment setting
C.1 Synthetic Experiment
We first generate directed graphs with 10 nodes and an edge density of 0.6, then randomly drop 60% of the edges. From each resulting subgraph, we randomly select two nodes and retain the subgraph if the path length between them is 3. In total, we collect 100,000 such graph cases. For training, we use a 5-layer decoder-only transformer with a hidden size of 96 and 12 attention heads. The model is trained for 20,000 iterations to ensure stable convergence.
For circuit tracing, we employ the ReLU version of the transcoder, with the dead feature window set to 50 and the L1 coefficient set to 0.0005. During visualization, we set the node ratio to 0.48 and the edge ratio to 0.86.
C.2 Real-world Experiment
The experimental settings for real-world data are summarized in Table 2. We include both the thinking prompt “Let’s think step by step” and the “<think>” tag. For the transcoders, we select parameter configurations that converge to at least an MSE value of 0.001 at the first layer.
Qwen3-0.6B Qwen2.5-1.5B Llama-8B DS-Qwen DS-Llama
Max Length MAWPS (1772) 200 200 200 400 400
GSM8K (1319) 300 300 300 500 500
ProntoQA (1500) 900 900 900 900 900
BoolQA (3270) 500 500 500 500 500
ARC (2376) 300 300 300 500 500
QASC (926) 300 300 300 500 500
Thinking prompt Math Let’s… Let’s… Let’s… Let’s… <> Let’s…<>
Logic Let’s… <> Let’s… Let’s… <> Let’s… <> Let’s… <>
Science Let’s… <> Let’s… Let’s… <> Let’s…<> Let’s… <>
Transcoder Edge Ratio 0.4 0.4 0.8 0.4 0.9
Node Ratio 0.5 0.5 0.9 0.5 0.9
L1_coefficient 0.0005 0.0005 0.00005 0.0005 0.00005
Training Tokens MAWPS 5*10^6 5*10^6 5*10^6 5*10^6 5*10^6
GSM8K 2*10^7 2*10^7 2*10^7 2*10^7 2*10^7
ProntoQA 1*10^7 1*10^7 1*10^7 1*10^7 1*10^7
BoolQA 5*10^6 5*10^6 5*10^6 2*10^7 2*10^7
ARC 2*10^7 2*10^7 2*10^7 2*10^7 2*10^7
QASC 2*10^7 2*10^7 2*10^7 2*10^7 2*10^7
Table 2: The experiment setting for real world cases. “Let’s…” denotes the prompt “Let’s think step-by-step”, and “<>” denotes “<think>”
Appendix D GraphGhost Results
In this section, we present the specific tokens ranked by graph algorithms. For the semantic merging case, where tokens are sorted by in-degree, Table 3 and Table 4 list the top tokens across different datasets and models. Table 7 highlights significant tokens that are merged across multiple layers of LLMs. For logic shifting, we report the top tokens for models and datasets in Table 5 and Table 6, respectively.
Top tokens Dataset
MAWPS GSM8K ARC QASC ProntoQA BoolQA
1 . the the , . the
2 the . , the uses ,
[…]
4 0 , . . ump is
5 : 2 are are is a
[…]
Table 3: The top tokens sort by the in-degree values across the datasets
Top token Model
Qwen3-0.6B Qwen2.5-1.5B DS-Qwen-1.5B Llama-8B-Instruct DS-Llama-8B
1 . <space> <space> <space> contin
2 the . , . <space>
3 <space> the . the **
4 0 = the , .
[…]
8 is has is 0 **
9 , had = of the
10 $n \n$ 0 and \\
Table 4: The top tokens sort by the in-degree values across the models
Top nodes Qwen3-0.6B Qwen2.5-1.5B DS-Qwen-1.5B Llama-8B-Instruct DS-Llama-8B
[…]
2 \_27 :_27 the_26 $\n\n_31$
[…]
4 the_26 :_26 ,_26 0_31 \**_31
5 :_27 \_27 ,_27 the_31 ._31
[…]
7 the_22 0_27 **_27 ,_31 **_31
[…]
9 0_26 \._27 ,_25 $\n\n_31$
[…]
11 of_26 :_25 to_27 $\n_30$
12 0_23 the_26 a_27 of_31 \_31
13 ._22 ._26 ._25 :_30 vase_0
14 :_24 \_26 the_22 is_31 ,_31
[…]
18 :_22 \._26 the_21 to_31 contin_29
[…]
Table 5: Top tokens ranked by PageRank across LLMs
Top tokens MAWPS GSM8K BoolQA ProntoQA SCQA ARC
1 0_27 the_27 the_27 ._26 the_27 the_27
2 \_27 0_27 the_26 is_27 ._27 a_27
3 ._27 ._27 $\n_27$ ,_27 ._27
4 the_26 \_27 a_26 ._25 ,_26 the_26
5 :_27 the_26 a_27 ,_25 are_27 a_26
6 :_26 :_27 ._27 ,_26 is_27 the_22
7 the_22 **_27 the_22 1_27 are_26 the_25
[…]
9 0_26 ,_27 the_23 ump_27 a_26 ,_27
10 ._26 ._24 ,_26 1_26 the_25 the_21
11 of_26 to_26 the_21 ._22 )_27 is_27
12 0_23 the_25 is_27 to_25 are_22 the_23
13 ._22 ,_26 ,_22 -_26 ._25 a_25
14 :_24 ._26 ,_21 given_23 the_22 a_22
15 :_25 (_27 ._24 the_26 )_26 a_24
[…]
18 :_22 the_22 a_23 \_25 to_26 and_27
19 ._23 to_25 ,_25 ster_27 are_24 ._24
20 the_23 ,_25 a_22 the_22 of_27 the_24
Table 6: Top tokens ranked by PageRank across datasets
Model Dataset Select Token Layers # In-degree New tokens Description
Qwen3-0.6B ARC the 1 26 is, not, of, in, and Basic function words
15 141 Because, But, could, bird Logical connectors/conditionals + noun
[…]
25 481 -check, A, 2), Carboon answer-related symbols and entities
MAWPS the 1 None neurons
15 40 Calculate, add, in Math calculation actions
20 100 1, \$, =, Answer, Step Answer template with math symbols
[…]
15 14 Question, Answer, Step Question–answer framing tokens
20 134 \$, 2, + divide Mathematical symbols and operations
25 147 How many, spent, together Math word-problem phrasing tokens
, 1 3 First, Next Sequencing/transition words
15 1 , Basic punctuation
20 11 Since, To Ttransition words
25 30 How, total Question words focusing on totals or amounts
Qwen2.5-1.5B MAWPS the 1 32 Calculate, add, of, in Function words + common math verbs
15 None neurons
20 38 Multiply, Question Math action words
25 54 1, Answer, step Answer-step related tokens
[…]
15 1 . Basic punctuation
20 49 =,+, got, Answer, Question Math operators mixed with answer framing tokens
25 63 \$, * Mathematical symbols
Llama-8BI MAWPS the 1 7 Question, For Question framing / introductory tokens
2 18 of, to, ing Basic function words
[…]
31 313 =, add, or Mathematical operators and logical connectors
. 1 56 1, yen, Question, to, ing Numeric units and question words
2 72 all, and, company Conjunctions and basic nouns
[…]
31 574 =+, 111, dogs Numbers combined with operators and concrete nouns
Table 7: The examples of how LLMs merge the semantics.

### v2

@@
 Abstract
-Large Language Models (LLMs) demonstrate remarkable reasoning capabilities, yet the structural mechanisms underlying these abilities remain under explored.
-In this work, we introduce GraphGhost, a unified framework that represents neuron activations and their signal propagation as graphs, explaining how LLMs capture structural semantics from sequential inputs and generate outputs through structurally consistent mechanisms.
-This graph-based perspective enables us to employ graph algorithms such as PageRank to characterize the properties of LLMs, revealing both shared and model-specific reasoning behaviors across diverse datasets.
-We further identify the activated neurons within GraphGhost and evaluate them through structural interventions, showing that edits to key neuron nodes can trigger reasoning collapse, altering both logical flow and semantic understanding.
-Together, these contributions position GraphGhost as a powerful tool for analyzing, intervening in, and ultimately understanding the structural foundations of reasoning in LLMs.
+Large Language Models (LLMs) exhibit strong reasoning capabilities on structured tasks, yet the internal mechanisms underlying such behaviors remain poorly understood.
+Existing interpretation methods mainly focus on token-level attributions, which provide limited insight into multi-step reasoning inside the model.
+We propose GraphGhost, a graph-based framework that models internal token interactions and neuron activations in LLMs as graphs.
+By aggregating token dependencies traced across layers, GraphGhost captures global information flow underlying model predictions.
+We formalize GraphGhost from two complementary perspectives: a sample view, which traces token dependencies for individual predictions, and a dataset view, which aggregates recurring structural patterns learned during training.
+Through graph analytics and quantitative experiments, we show that graph structural properties are closely associated with influential tokens and neuron nodes, and that perturbations to structurally critical nodes lead to measurable changes in reasoning behavior.
+These results indicate that the structural patterns captured by GraphGhost reflect meaningful internal organization of LLM reasoning.
+The codes are available at software part.
+Artifacts will be made available for research use only.
 GraphGhost: Tracing Structures Behind Large Language Models
-Xinnan Dai1, Kai Guo1thanks: Correspondence to guokai1@msu.edu.
-The code and data are available at https://github.com/DDigimon/GraphGhost, Chung-Hsiang Lo2, Shenglai Zeng1, Jiayuan Ding4, Dongsheng Luo3,Subhabrata Mukherjee4, Jiliang Tang1 1Michigan State University  2 Northeastern University 3 Florida International University  4 Hippocratic AI {daixinna, guokai1}@msu.edu
+Xinnan Dai1, XianXuan Long1, Chung-Hsiang Lo2, Kai Guo1thanks: Correspondence to guokai1@msu.edu.
+The code and data are available at https://github.com/DDigimon/GraphGhost, Shenglai Zeng1, Dongsheng Luo3, Jiliang Tang1 1Michigan State University  2 Northeastern University 3 Florida International University {daixinna, guokai1}@msu.edu
 1 Introduction
-Large Language Models (LLMs) have demonstrated promising capabilities in capturing structural information from text Sui et al.
+Large Language Models (LLMs) have demonstrated remarkable capabilities in solving structured reasoning tasks, where solutions require the composition of multiple intermediate steps governed by explicit relational or logical dependencies, such as mathematical problem solving, logical inference, and graph-based reasoning (Bommasani et al., 2022; Pan et al., 2023).
+Despite this success, the internal mechanisms by which LLMs represent, compose, and reuse such structural information during reasoning remain poorly understood (Nanda et al., 2023; Nananukul et al., 2025; Pandey et al., 2025).
+This is important because it provides a principled way to understand the source of intelligence in LLMs.
+Figure 1: From local explanations and human-aligned reasoning to global internal structures.
+(a) Token-level attribution explanations provide localized views of individual predictions.
+(b) Human-aligned step-by-step reasoning produces coherent reasoning traces without revealing internal mechanisms.
+(c) GraphGhost captures global token interactions inside LLMs as a structured graph, enabling mechanistic analysis of internal reasoning processes.
+Most existing efforts to understand LLMs’ behaviors on structured reasoning tasks focus on token-level explanations (Chen et al., 2025; Tak et al., 2025), as illustrated in Figure 1(a), which trace individual predictions back to contributing tokens or activations and provide only localized views of isolated outputs.
+Given that reasoning capability plays a central role in LLM performance across reasoning tasks (Wei et al., 2022; Lewkowycz et al., 2022) as shown in Figure 1 (b), existing token-level approaches provide limited insights into how representations are composed and reused across multiple reasoning steps (Geiger et al., 2025; Dziri et al., 2023).
+This limitation is increasingly recognized as a major obstacle to meaningful interpretability, as faithful understanding requires alignment between internal reasoning mechanisms and human-interpretable reasoning logic (Zhao et al., 2024; Turpin et al., 2023; Barez et al., 2025).
+To bridge this gap, we introduce GraphGhost, a graph-based framework that captures internal token interactions and neuron activations in LLMs as graphs as shown in Figure 1(c).
+Motivated by the central role of step-by-step reasoning in LLM performance, GraphGhost moves beyond isolated token attributions to explore how the reasoning steps are driven under the individual token explanation.
+This structural representation enables us to analyze reasoning as an internal, mechanistic process, revealing how intermediate tokens are merged, propagated, and reused during generation.
+Specifically, in Section 3, we formalize GraphGhost from two complementary perspectives.
+First, in the sample view, GraphGhost traces internal token dependencies within individual input–output pairs that contribute to a specific prediction, exposing reasoning structures for LLMs.
+Second, in the dataset view, GraphGhost is constructed by aggregating structures across samples to reveal recurring and shared token-level patterns learned from data.
+Through graph analytics, we demonstrate that LLM reasoning behavior is tightly coupled to internal graph structural properties, with influential tokens and neuron nodes characterized by structurally critical roles.
+These findings suggest that internal graph structures fundamentally shape how individual tokens influence LLM reasoning, through their structural roles within the model.
+In Section 4, we conduct extensive quantitative experiments to assess whether the structural patterns captured by internal graphs are meaningfully tied to LLM reasoning behavior.
+Using fidelity, infidelity, and sparsity metrics, we demonstrate that the tokens and neuron nodes identified by GraphGhost exert a significantly strong influence on model behavior.
+Furthermore, targeted interventions on structurally important nodes lead to measurable changes in reasoning trajectories and outputs, providing empirical evidence that GraphGhost captures meaningful internal structures governing LLM reasoning.
+In conclusion, we summarize our key contributions as follows:
+1) We propose GraphGhost, a unified framework for understanding how LLMs perform reasoning by modeling internal token interactions and neuron activations at both the sample and dataset levels.
+2) Using GraphGhost, we reveal that internal graphs capture structural constraints that govern model behavior, shaping information composition and propagation during reasoning.
+3) We conduct quantitative experiments to test whether graphs constructed by GraphGhost capture meaningful structural patterns of LLM reasoning.
+Our results show that structurally important token-related neuron nodes strongly influence reasoning trajectories, and that intervening on these nodes alters both language expression and logical flow.
+2 Related Work
+While Chain-of-Thought (CoT) prompting enhances LLM reasoning (Wei et al., 2022; Kojima et al., 2022), its faithfulness to the model’s intrinsic computations remains contested (Jacovi and Goldberg, 2020; Lyu et al., 2023).
+Prior work shows that CoT often acts as post-hoc rationalization, generating explanations that diverges from the underlying causal logic (Turpin et al., 2023) or internal attention patterns (Lanham et al., 2023; Barez et al., 2025).
+Moreover, such behavioral analyses provide limited insight into the specific neurons or circuits driving model predictions.
+To overcome the lack of causal rigor in probing Alain and Bengio (2018); Belinkov (2021) and attention analysis Clark et al.
+(2019); Vig and Belinkov (2019), research has pivoted to attribution methods (Sundararajan et al., 2017; Meng et al., 2023) and circuit tracing (Elhage et al., 2021).
+Recent methods leverages Sparse Autoencoders (SAEs) Bricken et al.
+(2023); Cunningham et al.
+(2023) and Transcoders Dunefsky et al.
+(2024) to decompose dense representations into interpretable features (Bloom et al., 2024; Ferrando et al., 2025; Paulo et al., 2025), enabling attributed graphs (Hanna et al., 2025; Marks et al., 2025) to model causal flow.
+GraphGhost extends this framework to uncover internal structures that affect LLM behavior in both sample and dataset view.
+Figure 2: We use a simple example: 2 + 2 * 4.
+(a) An attributed graph interpreting a single-token prediction.
+(b) GraphGhost (sample view) aggregates token dependencies along the reasoning chain.
+(c) GraphGhost (dataset view) captures cross-layer token correlations learned from the corpus.
+Here, “2_1” denotes token 2 at layer 1; “-_2” does not appear in the sample but is activated from other corpus contexts.
+Token “2_1” participates in both “2*4” and “2+8” computations.
+3 Tracing the GraphGhosts
+The reasoning behavior of LLMs emerges through successive token generations.
+Accordingly, existing interpretation methods primarily operate at the token level, attributing individual predictions to specific tokens or activations and providing localized explanations, as illustrated in Figure 2(a) Ameisen et al.
+(2025).
+However, reasoning does not arise from isolated tokens alone, but from structured interactions among multiple tokens within a sample and from recurring patterns shared across samples.
+To capture these aspects, we introduce two complementary views: a sample view and a dataset view.
+The sample view models token interactions within a single reasoning instance and formalizes this instance-specific structure as a GraphGhost (Figure 2(b); Section 3.1).
+The dataset view aggregates such structures across samples to reveal reusable reasoning patterns and is formalized in Section 3.2 (Figure 2(c)).
+3.1 Sample View
+LLMs have demonstrated strong capabilities in chain-of-thought reasoning.
+However, whether these reasoning chains faithfully correspond to coherent internal reasoning mechanisms remains unclear.
+Rather than attempting to answer this question conclusively, we begin with a toy example to illustrate how circuit-tracing methods can be used to probe the internal computations underlying such reasoning behaviors.
+Toy Example on Graph Reasoning
+Since search paths on a graph naturally decompose reasoning into node-by-node transitions, we adopt the path-reasoning task as a toy example to illustrate how nodes are selected at each step of the reasoning process.
+Following prior work Saparov et al.
 (2025); Wang et al.
-(2025, 2023).
-Notable examples include identifying relationships between entities in context Zhang and Soh (2024), reasoning and planning through structured thinking processes Yao et al.
-(2023a); Besta et al.
-(2024); Wu et al.
-(2024), and recognizing logical connections across paragraphs Pan et al.
-(2023).
-Trained on large-scale corpora, recent studies suggest that LLMs can transfer reasoning processes by extracting semantics and generalizing learned reasoning strategies Muennighoff et al.
-(2025).
-This raises the question of whether LLMs internally construct reasoning structures and leverage them across diverse tasks.
-Figure 1: An illustration of GraphGhost based on Qwen3-0.6B.
-a) GraphGhost illustrates a structural mechanism within LLMs that operates consistently across datasets.
-L indicates the layer number.
-Sequential inputs activate subsets of neurons, which merge tokens in the intermediate layers to capture semantic meaning.
-These activated neurons then propagate signals to the logit layers to generate outputs.
-b) Muting the neuron node.
-Changing the structure of GraphGhost leads to corresponding changes in both language expression and logical reasoning.
-“等于” means “equal to” in Chinese.
-To understand the mechanisms of constructing reasoning structures from text, prior studies have often simplified reasoning tasks into toy examples such as graph-based reasoning to examine how LLMs trace reasoning paths within giving structures Sanford et al.
-(2024).
-The underlying hypothesis is that uncovering how decoder-only transformers construct such graphs may yield insights that generalize to real-world tasks Saparov et al.
-(2025).
-While these toy graph settings provide preliminary insights—highlighting mechanisms such as token merging Dai et al.
-(2025b) and edge memorization Wang et al.
-(2024), they fall short when applied to complex, real-world structured data.
-To bridge this gap, we introduce GraphGhost, a framework that provides a structural perspective on how LLMs across different architectures perform reasoning tasks in diverse domains.
-GraphGhost captures neurons and their signal propagation inside LLMs, revealing how neurons are activated across samples and datasets, enabling shared semantic understanding and supporting the construction of reasoning processes across contexts.
-As illustrated in Figure 1 a), GraphGhost represents a structural mechanism modeled within LLMs, where neuron activations capture semantic information and guide token outputs.
-In the sentence-reading stage, LLMs merge tokens to derive semantic meaning, thereby activating specific neurons associated with those tokens.
-These activated neurons then propagate signals to higher layers, ultimately influencing the logit neurons in the output layers.
-Notably, while certain neurons are activated for the given samples, many remain silent.
-These silent neurons are reserved for unseen inputs from different domains, enabling shared output behaviors across contexts.
-For example, the symbol ‘X’ may activate the ‘=’ symbol in mathematical equations, even if the current sample does not require it, leaving those neurons in a silent state until needed.
-We define neurons as nodes and the signal propagation among neurons as edges.
-This representation allows us to analyze LLMs from a graph perspective.
-For instance, by inspecting high in-degree nodes or applying PageRank, we can identify key nodes that control next-token prediction, as further elaborated in Section 3.
-Furthermore, we evaluate the effectiveness of these nodes through structural interventions, such as node removal (Figure 1 b)).
-Removing specific nodes can lead to semantic shifts, such as switching to another language, or logical changes.
-In addition, we conduct statistical analyses of these structural perturbations to quantify their effects, as presented in Section 4.
-In conclusion, our key contributions are summarized as:
-  1. 1.
-    We introduce GraphGhost, a unified framework represented in graphs, explaining how LLMs capture structural semantics from sequential inputs and generate outputs through a structurally consistent mechanism.
-  2. 2.
-    We provide a graph-based perspective for understanding LLMs and apply graph algorithms on GraphGhost to characterize the properties of LLMs, revealing both unique and shared behaviors across different models and datasets.
-  3. 3.
-    We identify the activated neurons within LLMs and evaluate them by editing the structure of GraphGhost.
-Our experiments show that key neuron nodes can trigger reasoning collapse, altering both the logical flow and semantic understanding.
-2 Related work and Preliminary
-GraphGhost is built on circuit tracing, which is a typical LLM interpretation tool.
-In this section, we first introduce the background on LLM interpretability and reasoning tasks in Section 2.1 and Section 2.2, respectively.
-To illustrate how LLMs construct internal structures that align with explicit reasoning texts, we design a toy experiment in Section 2.3 to illustrate the alignment of inner signal propagation with explicit structure descriptions.
-Building on this foundation, we provide the preliminary definition in Section 2.4.
-2.1 Interpretation of LLMs
-Interpreting how LLMs perform reasoning is essential for aligning their behavior with human understanding Zhao et al.
-(2024).
-Existing approaches fall into two categories, self-explanation Gu et al.
-(2024) and causal-mechanistic interpretation Bloom et al.
-(2024).
-The details are suggested in Appendix A.1
-The causal-mechanism approach focuses on explaining how each token is predicted Anthropic (2023); Ameisen et al.
-(2025).
-Based on learning the internal reasoning pathways, Lindsey et al.
-(2025) introduced attribution graphs to capture token flows that contribute to final predictions.
-Formally, an attribution graph is defined as $G_{A}=(\hat{V},\hat{E},A)$, where $\hat{V}$ denotes activated neurons, $A_{s\to t},;s,t\in\hat{V}$ represents the contribution from node $s$ to $t$ (computed via backward Jacobians), and $\hat{E}$ indicates whether an edge exists.
-The edge set is controlled by an edge ratio, which filters the top influential connections according to $A$.
-2.2 Reasoning Tasks
-Reasoning tasks including mathmatical Cobbe et al.
-(2021), scientific Khot et al.
-(2020), and logic reasoning Clark et al.
-(2019) are set to evaluate the ability in solving complex problems.
-The details are in Appendix C.2
-Graph reasoning represents a distinct type of reasoning task Dai et al.
-(2025a).
-In this setting, models are provided with textual descriptions of graph structures and asked to reason over them.
-Success on such tasks suggests that LLMs can leverage implicit structural representations to guide reasoning Wang et al.
-(2023).
-However, how these implicit structures are internally constructed and maintained within LLMs remains an open question.
-2.3 Path tracing in the decoder-only transformer
-Graph reasoning tasks, particularly path reasoning, offer a prototypical setting for understanding how structural information in sequential inputs can be mapped to outputs that describe those structures Wang et al.
-(2023).
-In this task, the model is provided with a graph description and a query consisting of a start node and an end node, and the objective for LLMs or decoder-only transformers is to predict the path connecting the two nodes Dai et al.
-(2024); Wang et al.
-(2024).
-We adopt this basic setup to investigate how decoder-only transformers perform path reasoning.
-Following prior work Wang et al.
-(2024), we use simplified graph descriptions, and a query (“from start point S to end point E”) and must predict the connecting path.
-as illustrated in Figure 2(a), and train a 5-layer GPT-2 architecture decoder-only transformer on this task.
-The trained model achieves 98% accuracy on the test set.
-We then apply the circuit-tracing method, specifically using transcoder Dunefsky et al.
-(2024), to analyze how the transformer models the path-finding process, as shown in Figure 2(b).
-Full experimental details are provided in Appendix C.1.
-Circuit tracing reveals how decoder-only transformers review the previous context to identify related nodes and predict the next token.
-From a structural understanding perspective, neighboring nodes are merged in the middle layers.
-For example, at the third layer, node 4 is merged with node 5, forming the semantic representation of the edge between them.
-Further, in generating the output tokens, the final token in the sequence receives the message that an edge exists between 4 and 5 at layer 3, leading the model to predict 5 as the next token.
-From this toy example, we observe that decoder-only transformers are not only capable of reasoning about paths within given graph structures, but that their underlying mechanism involves activated neurons propagating signals across layers.
-However, beyond the toy example with decoder-only transformers, does this mechanism extend to LLMs that can solve more complex tasks expressed in natural language?
-To address this, we introduce GraphGhost, a unified framework that represents these neuron activations and their signal flows as graphs.
-GraphGhost enables us to trace and mergeattribution graphs defined in Section 2.1 across diverse reasoning tasks, revealing how LLMs encode and reuse structural reasoning patterns.
-In the following sections, we apply GraphGhost to multiple models and domains to analyze how their internal reasoning structures build up, interact, and generalize.
+(2024) that studies LLM reasoning and planning through simplified graph abstractions, we focus on shortest-path search as our canonical setting, as illustrated in Figure 3(a).
+Graphs are specified as lists of edges, with the symbol “|” used to separate individual edges.
+In the queries, we denote the start and end nodes using the markers “S” and “E”, respectively.
+In the answers, we use the symbol “P” to indicate that the model is required to generate an explicit reasoning path.
+we train a 5-layer GPT-2 architecture decoder-only transformer on this task.
+The trained model achieves 98% accuracy on the test set.To understand how the trained model internally performs path finding, we analyze its inference process using circuit-tracing methods.
+Specifically, following Hanna et al.
+(2025), we train a transcoder Dunefsky et al.
+(2024) for each layer of the transformer.
+Each transcoder takes the layer-wise hidden activations as input and learns a sparse feature representation that reconstructs these activations, yielding token-level semantic features for each MLP layer.
+We then replace the original MLP layers with the corresponding transcoders and compute the logit-level contribution weights of each transcoder feature to the final prediction.
+Using transcoder features as nodes and their logit contribution weights as directed edges, we construct the attributed graph.
+The resulting attribution graph, visualized in Figure 3(b), illustrates how the transformer models the path-finding process.
+Full experimental details are provided in Appendix B.1.
+Circuit tracing reveals that decoder-only transformers revisit the preceding context to identify relevant nodes and predict the next token.
+From a structural perspective, neighboring nodes are progressively merged in the middle layers.
+For example, at the third layer, node 4 is merged with node 5, forming a semantic representation of the edge between them.
+During generation, the final token in the sequence receives the signal that an edge exists between nodes 4 and 5 at layer 3, which leads the model to predict node 5 as the next token.
+This mechanism differs from human graph reasoning, which explicitly retrieves relations, checks connectivity, and composes shortest paths.
+We then evaluate this assumption on real LLM reasoning tasks.
 (a) Graph description in a simplified way. The sequence includes the graph description, query, and answer.
@@
 (a) The simplified descriptions to predict the path between 4 and 7; (b) The circuit tracing for the interpretation of why token 5 is selected to be predicted.
-2.4 Preliminary definitions
-We formalize the core concepts underlying GraphGhost.
-GraphGhost is constructed from attribution graphs obtained using $\text{CircuitTracer}(\cdot)$, which traces neuron activations contributing to next-token prediction and records their signal flow.
-Let $X=(t_{1},\dots,t_{n})$ be the input tokens of an LLM with layers $\{f^{(1)},\dots,f^{(L)}\}$.
-For a target token $t_{n+1}$, the attribution graph is defined as $A_{n+1}=\text{CircuitTracer}(t_{n+1}),$ where each node represents an activated neuron $(t_{i},\ell)$ with activation $h^{(\ell)}(t_{i})=f^{(\ell)}(h^{(\ell-1)}(t_{i}))$, and each directed edge $w_{s\to t}$ encodes the contribution from neuron $s$ to $t$.
-The set of contributing tokens is $\hat{T}(t_{n+1})=\{\,t_{i}\mid(t_{i},\ell)\in A_{n+1}\,\}$.
-where $\ell$ denotes the layer index.
-These local attribution graphs form the building blocks of GraphGhost, which aggregates them across samples into a unified weighted structure.
-Following the definition of attribution graphs in Ameisen et al.
-(2025), a neuron node is represented as a pair $(t_{i},\ell)$, where $t_{i}$ denotes a token and $\ell$ denotes the layer index.
-Each neuron node has a corresponding hidden activation $h^{(\ell)}(t_{i})$, representing the hidden state of token $t_{i}$ at layer $\ell$, computed as $h^{(\ell)}(t_{i})=f^{(\ell)}\big(h^{(\ell-1)}(t_{i})\big).$ For a source neuron node $s=(t_{i},\ell_{s})$ and a target neuron node $t=(t_{j},\ell_{t})$ with $\ell_{s}<\ell_{t}$, the edge weight in the attribution graph is defined as $w_{s\to t}=A_{s\to t},$ where $A_{s\to t}$ denotes the attribution score from $s$ to $t$.
-For example, in Figure 2,to generate token $5$, the contributing tokens are $\{4,5,7,\texttt{"|"},\texttt{"S"},\texttt{"P"}\}$.
-The corresponding neuron nodes include $h^{(3)}(5)$, $h^{(4)}(7)$, and $h^{(5)}(4)$, among others.
-The edge weight from $h^{(3)}(5)$ to $h^{(5)}(4)$ is denoted as $w_{h^{(3)}(5)\to h^{(5)}(4)}$.
-3 GraphGhost in LLMs
-Beyond the toy examples, we apply GraphGhost to real-world LLMs to uncover their internal reasoning structures.
-GraphGhost traces neuron activations that contribute to next-token generation, represents them as local attribution graphs, and merges these across samples into a global weighted graph.
-This unified structure reveals stable signal propagation patterns shared across tasks and models.
-Section 3.1 details the tracing and aggregation process.
-To evaluate whether these structures correspond to meaningful reasoning behaviors, we introduce the experimental setup in Section 3.2.
-which specifies the models, datasets, and tracing configurations used for constructing GraphGhost.
-Section 3.2 and Section 3.4 then present case studies and quantitative analyses based on these settings.
-3.1 Tracing the GraphGhost
-To construct a unified GraphGhost that captures reasoning structures shared across multiple samples, we aggregate the attribution graphs obtained from each sample.
-Given a set of samples $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$, where each sample $s_{i}$ contains an answer token $t_{n}^{(i)}$, we initialize a global graph $G=(V,E,W)$ with empty node, edge, and weight sets.
-For each sample $s_{i}$, we start from its answer token $t_{n}^{(i)}$ and iteratively trace the contributing tokens and their local attribution graph using the $\text{CircuitTracer}(\cdot)$ function.
-Each traced graph produces a set of neuron nodes $\hat{V}$ and directed edges $\hat{E}$, which are merged into the global GraphGhost.
-If an edge $e\in\hat{E}$ does not exist in $E$, it is added with an initial weight $W(e)=1$; otherwise, its weight is incremented by one.
-This cumulative weighting reflects the frequency of edge occurrences across samples, capturing the stability and recurrence of signal propagation pathways.
-After processing all samples, the resulting GraphGhost $G=(V,E,W)$ represents a multi-sample structural abstraction of the model’s internal reasoning dynamics, which is a weighted directed graph describing neuron-level signal propagation.
-The full procedure is summarized in Algorithm 1 fromAppendix B.
-3.2 Experiments setting
-We evaluate a range of LLMs and datasets to investigate whether GraphGhost constitutes a general mechanism across models.
-Specifically, we consider Qwen3-0.6B (28 layers), Qwen2.5-1.5B (28 layers), Llama3.1-8B-Instruct (32 layers), DeepSeek-distilled-Qwen2.5B-1.5B (DS-Qwen-1.5B) (28 layers), and DeepSeek-distilled-Llama-8B (DS-Llama-8B) (32 layers).
-For datasets, we examine three domains: (i) mathematical reasoning (GSM8K, MAWPS), (ii) logical reasoning (ProntoQA, BoolQA), and (iii) scientific reasoning (ARC-Easy, QASC).
-To elicit the models’ reasoning processes, we apply Chain-of-Thought prompting.
-Following Hanna et al.
-(2025), we then train transcoders at every layer for each dataset–model pair to probe neuron activations during reasoning, ensuring that the transcoders effectively capture the most strongly activated neurons.
-All the experiments are run on a single H200 GPU.
-Full experimental details are provided in Appendix C.2.
-3.3 Case study on GraphGhost
-Figure 3: Case study of tracing contributing tokens for the answer 14 in Qwen3-0.6B model.
-We highlight the numerical reasoning flow captured by GraphGhost to visualize how number-related tokens contribute to the final prediction.
-We further provide some case studies of GraphGhost on contributing nodes and substructures, respectively.
-Figure 3 is an illustrative example of tracing contributing tokens within a single sample.
+GraphGhost in Sample Interpretation
+In the toy example, we observe that reasoning in basic decoder-only transformers is driven by token-level associations, where previously observed in-context tokens contribute to information aggregation during token generation, rather than through explicitly modeled logical structures.
+Motivated by this observation, we merge answer-related circuits to examine how reasoning is internally realized for a specific query in LLMs.
+At the sample level, GraphGhost aims to capture the internal structural dependencies that contribute to a model’s final prediction.
+Given a sample $s$ with final answer token $t_{n}$, we initialize an empty graph $G^{\text{sample}}=(V,E)$ together with a local edge-weight accumulator $W^{\text{local}}$, which is used to aggregate contribution strengths when the same dependency is discovered multiple times during backward expansion.
+The graph construction starts from $t_{n}$ and iteratively expands backward through the model’s internal computation.
+For each token $t$ in the current frontier that is not part of the question tokens, we apply a circuit tracing procedure to extract the local logit-based attribution graph rooted at $t$ follows the previous paragraph.
+This step yields a set of contributing tokens as sample-view nodes $\hat{V}$, directed edges $\hat{E}$, and local attribution weights $W^{\text{loc}}$, characterizing how information flows to the prediction of $t$.
+Newly discovered tokens are added to the frontier to enable recursive exploration of upstream dependencies.
+The nodes and edges from each local attribution graph are by taking the union over nodes and edges, removing duplicates and retaining only the existence of dependency relations.
+As a result, the sample-level GraphGhost is an unweighted structural graph that captures which token-level dependencies are present, independent of their individual contribution strengths.
+In parallel, local edge weights are accumulated in $W^{\text{local}}$ for each observed edge, which is retained solely for later aggregation at the dataset level.
+The final output is a sample-specific GraphGhost graph $G^{\text{sample}}=(V,E)$ together with the local edge-weight map $W^{\text{local}}$.
+The overall algorithm is shown in Algorithm 1 in Appendix A.
+Figure 4: Case study tracing tokens contributing to the answer 14 in Qwen3-0.6B, illustrating how GraphGhost captures the numerical reasoning flow linking number-related tokens to the final prediction.
+The sample view of GraphGhost characterizes step-by-step reasoning in LLMs by modeling structural dependencies among tokens as reasoning unfolds.
+Here, we show an example in Figure 4.
 The identified token set $T$ includes tokens from both the input question (highlighted in green) and those generated along the reasoning chain (in orange).
@@
 This tracing process demonstrates how local dependencies between tokens collectively form reasoning paths during sequential generation.
-Figure 4 demonstrates a representative subgraph extracted from the GraphGhost constructed on Qwen3-0.6B model, integrating neuron activations from mathematical reasoning (GSM8K and MAWPS), scientific reasoning (ARC), and logical reasoning (ProntoQA).
-Each neuron node is annotated with its corresponding token and layer ID.
-For example, the activated period at layer 22 is denoted as $_{22}.$ This figure highlights how neuron activations from diverse domains converge into a unified structural representation, revealing consistent signal propagation patterns underlying different types of reasoning tasks.
+3.2 Dataset View
+The reasoning capabilities of LLMs are learned from large-scale corpora, which motivates the assumption that certain neuron units within LLMs encode reusable token-level connections that support generalization to new reasoning tasks.
+Accordingly, we introduce the dataset-level view of GraphGhost to explore the implicit structures inside LLMs.
+3.2.1 Method
+To build the dataset-level of GraphGhost, we aggregate sample-level reasoning structures to reveal recurring internal information-flow patterns shared across multiple samples from datasets.
+Given a set of samples $\mathcal{S}=\{s_{1},\ldots,s_{m}\}$, we construct a unified graph for the dataset as $G^{\text{data}}=(V^{\text{data}},E^{\text{data}},W^{\text{data}})$.
+For each sample $s_{i}\in\mathcal{S}$, we first obtain its sample-level GraphGhost $G^{(i)}=(V^{(i)},E^{(i)})$ together with a local edge-weight map $W^{\text{local}}$ using the sample-view construction.
+The dataset-level node and edge sets are formed by merging sample graphs via set union: $V^{\text{data}}=\bigcup_{i}V^{(i)}$ and $E^{\text{data}}=\bigcup_{i}E^{(i)}$.
+In parallel, edge attribution strengths are accumulated across samples.
+For each edge $e\in E^{(i)}$, we update its dataset-level weight as $W^{\text{data}}(e)\leftarrow W^{\text{data}}(e)+W^{\text{local}}(e),$ where $W^{\text{local}}$ captures unnormalized, sample-specific contribution strengths.
+This accumulation preserves both the frequency and strength of recurring dependencies across samples.
+After aggregating all samples, we apply row-stochastic normalization to $W^{\text{data}}$ to obtain relative edge importance at the dataset level.
+The resulting graph captures recurring reasoning paths, while the normalized weights quantify their global importance across the dataset.
+The full procedure is summarized in Algorithm 2 in Appendix A.
+The dataset view aims to reveal whether LLMs exhibit recurring internal structures across different tasks and datasets.
+To this end, we visualize a subgraph extracted from the dataset-level GraphGhost in Figure 5, which aggregates sample-level GraphGhosts constructed from four datasets spanning different domains.
+In this graph, each node represents a token–layer pair identified by GraphGhost, and node colors indicate the datasets in which the token participates, while mixed colors denote tokens shared across multiple datasets.
+From this visualization, several consistent patterns emerge.
+First, certain tokens appear across all four datasets, suggesting shared internal mechanisms.
+For example, the period symbol “.” at layer 22 exhibits strong logit contributions across all datasets, and the discourse marker “So” frequently appears in diverse reasoning chains.
+Second, domain-specific patterns are also evident: arithmetic-related tokens such as “Adding” and expressions like “6+” are shared between GSM8K and MAWPS, which both involve numerical reasoning.
+In contrast, some nodes are task-specific; for example, $\mathrm{B_{18}}$ appears only in ARC, reflecting choice-based reasoning.
+This shows that dataset-level GraphGhost captures both shared reasoning primitives and task-specific structures.
 Figure 5: A subgraph of GraphGhost in Qwen3-0.6B model. We include 4 datasets from different domains.
-In the subgraph, we observe that some nodes are activated consistently across datasets, while others contribute only to specific domains or even individual samples.
-For example, the period symbol “.” at layer 22 triggers strong logit activations across all four datasets, and the token “So” frequently appears across different reasoning types.
-Domain-specific reasoning tokens also emerge within the same domain.
-For instance, the action “Adding” and arithmetic expressions such as “6+” appear in both GSM8K and MAWPS.
-In contrast, certain nodes are unique to individual samples—for example, $\mathrm{B_{18}}$ activates only in choice-option reasoning within the ARC dataset.
-3.4 Analysis on GraphGhost
-In Figure 4, we also observe two typical actions associated with the node $_{22}.$ The first is semantic understanding, exemplified by the “6+” operation.
-The second is logic shifting, where the activation leads to the token “So” in the subsequent prediction.
-We further analyze these two actions from a statistical perspective.
-Semantic understanding
-In Section 2.3, we discussed how edge information is merged within the inner layers.
-Building on this, we analyze the token-merging behavior of neuron nodes by examining their average in-degree across layers in LLMs’ settings.
-The statistical results, summarized in the Appendix D and illustrated in Figure 5, show the top tokens across different datasets and models.
-We observe that the tokens with the highest in-degree are domain-sensitive.
-For example, in the ProntoQA dataset, the model is more sensitive to the period token than to the comma.
-This suggests that LLMs activate specific neurons in response to domain-specific data to capture semantics.
-Across models, however, the activated neurons are generally consistent: tokens such as <space>, period, and “the" consistently appear among the top-ranked, and this alignment persists even after finetuning.
-Figure 5: The top merging nodes across the models and datasets.
-We further analyze the merging behavior across layers by computing the in-degree ratio.
+3.2.2 Analysis
+Beyond visualization, we conduct further graph-based analyses to examine how the structures identified by GraphGhost correspond to LLM behaviors.
+Specifically, we analyze the structures from two perspectives: (1) token-merging behavior, quantified by node in-degree; and (2) trigger tokens that initiate reasoning, identified using PageRank.
+In addition, we present case studies to illustrate how perturbations based on GraphGhost-identified structures affect the progression of the reasoning process.
+We consider Qwen3-0.6B (28 layers), Qwen2.5-1.5B (28 layers), Llama3.1-8B-Instruct (32 layers), DeepSeek-distilled-Qwen2.5B-1.5B (DS-Qwen-1.5B; 28 layers), and DeepSeek-distilled-Llama-8B (DS-Llama-8B; 32 layers).
+We evaluate these models on three reasoning domains: (i) mathematical reasoning (GSM8K, MAWPS), (ii) logical reasoning (ProntoQA, BoolQA), and (iii) scientific reasoning (ARC-Easy, QASC).
+We obtain the GraphGhosts for each model and dataset.
+All experiments are conducted on a single H200 GPU.
+Each sample will run 0.5-3h, depended on the length of reasoning tokens.
+Full experimental details are provided in Appendix B.2.
+Semantic Merging
+We begin by analyzing semantic merging behavior using the average in-degree of neuron nodes across layers.
+We assume that multiple tokens converge within the network, indicating potential semantic consolidation.
+For instance, in the representation of the number “10”, the tokens “1” and “0” exhibit convergent interactions in intermediate layers, consistent with the formation of a unified numerical representation.
+In the GraphGhost representation, in-degree quantifies how many contributing nodes are merged into a target node.
+Accordingly, nodes with higher in-degree correspond to tokens that integrate information from multiple sources and serve as central aggregation points for reasoning-related signals.
+Based on this, we examine which layers and tokens are preferentially in LLMs by GraphGhost structures.
+The statistical results are summarized in Appendix C and visualized in Figure 6.
+We observe that tokens with high in-degree exhibit clear domain sensitivity.
+For example, in the ProntoQA dataset, the model is more sensitive to the period token than to the comma, suggesting that punctuation plays different semantic roles across reasoning domains.
+Across models, however, the dominant tokens are largely consistent: special tokens such as <space>, the period, and “the” consistently rank among the top, and this pattern persists even after finetuning.
+Figure 6: The top merging nodes evaluated by in-degree ratio. Left: The highest score of the in-degree ratio across datasets. Right: The highest score of the in-degree ratio across the models
+We further set the in-degree ratio to measure the merging behavior across the datasets and models.
 Let $d_{i}^{(\ell)}$ denote the in-degree of neuron node $(t_{i},\ell)$, where $\ell$ is the layer index and $i$ indexes the neuron within that layer.
 The in-degree ratio of layer $\ell$ is then defined as $R_{\ell}=\frac{\sum_{i=1}^{N_{\ell}}d_{i}^{(\ell)}}{\sum_{k=1}^{L}\sum_{i=1}^{N_{k}}d_{i}^{(k)}},$ where $N_{\ell}$ is the number of neurons in layer $\ell$ and $L$ is the total number of layers.
+A higher in-degree ratio $R_{\ell}$ implies that layer $\ell$ concentrates a larger fraction of merging behavior, indicating that many reasoning-related signals converge at this layer.
+In contrast, a lower ratio suggests that neurons in the layer are less involved in integrating information, playing a more distributive or transitional role in the reasoning process.
 As shown in Figure 7, the in-degree ratio distribution is imbalanced across both models and tokens.
@@
 Figure 7: The in-degree proportions on various layers with MAWPS dataset
-Logic shifting.
-The reasoning structure is segmented by special tokens such as So, Step, and Wait Muennighoff et al.
-(2025); Bogdan et al.
-(2025).
-We assume that activated neuron nodes in the middle layers trigger these special token expressions at the logit layers.
-At the same time, the neuron nodes themselves are activated by input tokens.
-To capture both effects, we treat logit-layer nodes as signals directed back to the neuron nodes by reversing their edge directions.
-We then apply the PageRank algorithm to identify the most influential nodes.
-We summarize the results in Figure 7 and Figure 8, and the full details can be found in Appendix D.
-Figure 7: Top contributing tokens and layers in Qwen3-0.6B in various datasets.
-Figure 8: Top contributing tokens and layers across models evaluated by MAWPS dataset.
-Figure 7 presents the token distribution across datasets, highlighting the distinction between silent and activated tokens in LLMs when processing data from different domains.
-For example, numerical tokens such as 0 are more frequently activated in mathematical datasets, while general tokens like the period and “the" appear across multiple domains.
-Furthermore, Figure 8 illustrates that different models exhibit distinct activation preferences.
-After finetuning with distilled data, DS-Qwen-1.5B and DS-Llama-8B show increased activation of the token "**", suggesting that DeepSeek-distilled data encourages reliance on this token during generation.
-From a layer-level perspective, activated nodes consistently emerge in higher layers across both models and datasets, though not exclusively in the final layer; in the Qwen series, layers 22 to 27 frequently contain critical activated nodes.
-4 GraphGhost affects the reasoning process
-GraphGhost paves a new way in which LLM reasoning processes can be analyzed and controlled.
-To evaluate this, we perform perturbation experiments at two complementary levels.
-(i) At the input level, we modify the contributing tokens identified in Section 4.1 to examine how altering key linguistic elements affects the model’s reasoning trajectory and output behaviors.
-(ii) At the neuron level, we intervene on the influential neuron nodes detected via in-degree analysis and PageRank in Section 4.2, assessing how structural disruptions within the reasoning pathways influence the model’s internal logic flow.
-Together, these two perturbation perspectives reveal how surface-level token changes and internal structural shifts jointly shape the reasoning dynamics captured by GraphGhost.
-4.1 Impact of contribution tokens
-GraphGhost in LLMs leverages contributing tokens to capture context and perform reasoning.
-In practice, we assess their importance by masking tokens, replacing each with a blank token.
-To quantify this effect, we evaluate two metrics: Sparsity and Fidelity.
-Sparsity measures the proportion of contributing tokens in a given question, while Fidelity evaluates whether the answer generated from only the contributing tokens remains consistent with the original answer.
-Formally, $\mathrm{Sparsity}=\frac{\#C}{\#T}$ and $\mathrm{Fidelity}_{T}=\mathds{1}(LLM(C),LLM(T))$, where $C$ denotes the set of contributing tokens, $T$ is the full set of tokens, and $\mathds{1}$ is an indicator function returning 1 if the two answers are identical and 0 otherwise.
-In practice, we use gemini-2.5-flash-lite to identify whether the two predictions are the same, with the prompt as: "Comparing two answers for a question: Answer1 {LLM(T^)LLM($\hat${T})}, Answer2 {LLM(T^)LLM($\hat${T})}.
-Do they have the same final results?" To compare the effectiveness of GraphGhost across models, we evaluate 100 samples that are solvable by Qwen3-0.6B, Qwen2.5-1.5B, and DS-Qwen-1.5B.
-The results are presented in Figure 9.
-Figure 9: Sparsity Vs. Fidelity in various datasets and models.
-Change the semantic Change the language Question: A ) nervous and digestive B ) nervous and circulatory C ) respiratory and digestive D ) respiratory and circulatory Question: Which two body systems would most directly remove extra fluid from a person’s lungs?
+Figure 8: Top contributing tokens and layers in the dataset-level GraphGhost of Qwen3-0.6B across different datasets.
+Left: The most frequently contributing tokens.
+Right: The most frequently contributing tokens.
+Figure 9: Top contributing tokens and layers across models evaluated by MAWPS dataset. Left: The most frequently contributing tokens. Right: The most frequently contributing tokens.
+Case studies Semantic merging controls language Question: A ) nervous and digestive B ) nervous and circulatory C ) respiratory and digestive D ) respiratory and circulatory Question: Which two body systems would most directly remove extra fluid from a person’s lungs?
 Original answer: … B) Nervous and circulatory systems: The nervous system is responsible for regulating the body’s response to stress, while the circulatory system is responsible for transporting blood throughout the body.
 Neither of these systems directly remove extra fluid from the lungs… Mute the neuron node ’the’ at layer 20 Perturbed answer: … B) Nervous and circulatory systems: The nervous system is responsible for controlling the - circulatory system is responsible for transporting blood throughout the分区肺.
-While the分区肺需要血液来清除多余的水分,这些系统本身并不直接参与这个过程。…   Change the word expression Question: To determine how many trips Jerry will make, we need to calculate the total number of trays he needs to pick up and then divide by the number of trays he can carry at a time.
-Original answer: … we divide the total number of trays, 16.0 trays ÷ 8.0 trays per trip = 2 trips… Mute the neuron node ’.’ at layer 17 Perturbed answer: … Number of trips = Total trays / 8 = 16 / 8 = 2 trips …
-Figure 10: Changing the semantic understanding. The characters in changing the language case is translating the original answer.
-As shown in Figure 9, compared with random masking of the given question, the contributing tokens identified by GraphGhost successfully preserve the generated answer.
-This effect is particularly significant in mathematical reasoning tasks such as MAWPS and GSM8K, where even when half of the question tokens are removed, LLMs are still able to produce the correct prediction.
-In contrast, the performance gap between randomly masked and selected tokens is smaller in scientific reasoning tasks such as ARC dataset.
-This is because in ARC the answers often involve selecting the correct option, and in some cases LLMs rely on direct recognition of the options from the given context rather than explicit reasoning.
-4.2 Modification on the graph structure
-To evaluate whether GraphGhost successfully identifies the most influential tokens in LLMs, we mute specific neuron nodes and observe the resulting changes in the outputs.
-Specifically, neuron nodes are muted by setting the corresponding positions of their layer–token pairs to zero during the reasoning process.
-We evaluate the impact through both case studies and statistical analysis.
-Case study
-First, as discussed in Section 3.4, the impacts include changes in semantic understanding and logic shifting.
-As shown in Figure 10, the semantic understanding is reflected in changing the language formulation and word expression.
-When the neuron node “the” is deleted at layer 20, the English semantic pathway is disrupted, activating Chinese neurons that share the same semantic signal.
-As a result, the output appears in Chinese, directly translated from English.
-Furthermore, in the word-expression change case, deleting the neuron “.” at layer 17 causes the token “/” to replace “÷.” This suggests that another neuron is responsible for propagating the division signal, expressing it as “/.” These findings indicate that LLMs store semantics redundantly across multiple neuron nodes: when the most relevant neuron is muted, the model relies on alternative neurons that produce similar expressions.
-In Figure 11, we demonstrate how GraphGhost influences reasoning by altering key steps or driving the generation into loops.
-In the key-step change case, deleting the neuron node “.” at layer 25 introduces a new reasoning token such as “So", ultimately leading the model to produce the correct answer.
-In the loop case, deleting the neuron node “the” at layer 27 causes the reasoning to degenerate into a repetition, producing “theShaper Information.” These cases suggest that even a single neuron in the middle layer can fundamentally change the outcome of the reasoning process.
-Change the reasoning structure Change the key steps Question: There are 7.0 crayons in the drawer and 6.0 crayons on the desk.
+While the分区肺需要血液来清除多余的水分,这些系统本身并不直接参与这个过程。…   Reasoning trigger changes logic Question: There are 7.0 crayons in the drawer and 6.0 crayons on the desk.
 Sam placed 4.0 crayons and 8.0 scissors on the desk.
@@
 … - So, the total number of crayons = 7.0 (drawer) + 10.0 (desk) = 17.0 crayons.
-Failed into loop Question: Alyssa picked 25.0 limes and Mike picked 32.0 limes.
-Tom picked 12.0 plums.
-How many limes were picked in all?
-Let’s think step by step Original answer: Step 1: Identify the number of limes picked by Alyssa and Mike.
-Alyssa picked 25.0 limes.
-Mike picked 32.0 limes.
-… Mute the neuron node ’the’ at layer 27 Perturbed answer: Step 1: Identify theSharper Information In this problem, we are given theSharper Information about theSharper Information about Alyssa, Mike, and Tom’s limes and plums.
-We need to find theSharper Information about theSharper Information about theSharper …
-Figure 11: Changing on the logic shifting
-Statistical analysis
+Figure 10: Semantic control and reasoning triggers change the model generation processing. In the language-switching case, the generated characters are direct translations of the original answer.
+Reasoning Triggers
+The reasoning structure is segmented by special tokens such as So, Step, and Wait Bogdan et al.
+(2025).
+We assume that activated neuron nodes in the middle layers trigger these special token expressions at the logit layers.
+At the same time, the neuron nodes themselves are activated by input tokens.
+To capture both effects, we treat logit-layer nodes as signals directed back to the neuron nodes by reversing their edge directions.
+We then apply the PageRank algorithm to identify the most influential nodes.
+We summarize the results in Figure 8 and Figure 9, and the full details can be found in Appendix C.
+Figure 8 presents the token distribution across datasets, highlighting the distinction between silent and activated tokens in LLMs when processing data from different domains.
+For example, numerical tokens such as 0 are more frequently activated in mathematical datasets, while general tokens like the period and “the" appear across multiple domains.
+Furthermore, Figure 9 illustrates that different models exhibit distinct activation preferences.
+After finetuning with distilled data, DS-Qwen-1.5B and DS-Llama-8B show increased activation of the token "**", suggesting that DeepSeek-distilled data encourages reliance on this token during generation.
+From a layer-level perspective, activated nodes consistently emerge in higher layers across both models and datasets, though not exclusively in the final layer; in the Qwen series, layers 22 to 27 frequently contain critical activated nodes.
+Case Study
+We further conduct case studies to illustrate how semantic merging and reasoning-trigger effects influence the reasoning process.
+Based on the analyses in the previous sections, we observe that tokens, such as “the” and “.”, are identified as important by both analysis methods, although their functional roles vary across layers.
+For instance, in Qwen3, these tokens tend to merge in middle layers (e.g., around layer 15), while in higher layers (e.g., beyond layer 25), they often act as reasoning triggers.
+Accordingly, we design case studies to examine how the roles of such tokens evolve across layers during reasoning.
+Specifically, we mute neurons in the LLM when they correspond to a given token at a specified layer.
+We show that muting neurons in middle layers leads to semantic changes, as illustrated in Figure 10.
+When the neuron node associated with “the” is deleted at layer 20, the English semantic pathway is disrupted, activating Chinese neurons that share similar semantic signals.
+As a result, the model produces output in Chinese, effectively translating from the original English text.
+In the Reasoning trigger case, deleting the neuron node “.” at layer 25 introduces a new reasoning token such as “So", ultimately leading the model to produce the correct answer.
+The full results are shown in Appendix B.3.
+Figure 11: Sparsity Vs. Fidelity in various datasets and models.
+4 Quantitative Experiments
+To evaluate whether the structural patterns identified within LLMs are functionally relevant to reasoning behavior, We perform quantitative experiments that examine whether GraphGhost-identified tokens exhibit consistent and measurable significance.
+We first introduce the evaluation metrics in Section 4.1.
+The evaluation considers both the sample-level view (Section 4.2) and the dataset-level view (Section 4.3).
+4.1 Metrics
 […]
-We first evaluate the tokens ranked by the PageRank algorithm.
-Specifically, we calculate the average InFidelity score when muting the corresponding neuron nodes and identify the maximum-impact neuron nodes within the top-$k$ positions.
-The results, summarized in Table 1, include the top-1–10 and top-11–20 nodes.
-Overall, the top-10 nodes show a markedly higher influence on the model’s reasoning process, indicating that the PageRank algorithm effectively identifies influential nodes in the GraphGhost framework.
-Moreover, larger models demonstrate greater stability than smaller ones—for instance, Llama-8B achieves an InFidelity score of 0.28, whereas Qwen3 reaches 0.58.
-Interestingly, the <space> token emerges as the most impactful token, exerting a strong influence on overall model performance, particularly in higher layers.
+4.2 Contribution Token Selection in GraphGhost: Sample View
+To evaluate whether GraphGhost can identify tokens that are functionally important for the reasoning process, we compare it against attention-based explanation methods (Clark et al., 2019; Liu et al., 2024) and random baselines.
+Instead of analyzing individual tokens, we adapt the same pipeline as GraphGhost to evaluate its effectiveness.
+The detailed settings are provided in Appendix B.4.
+The comparative results are summarized in Figure 11.
+For a detailed breakdown of the performance for each individual model without averaging, we provide the complete results in Appendix D (Figure 15).
+GraphGhost consistently outperforms both random and attention-based baselines.
+This advantage is most pronounced on mathematical reasoning tasks (MAWPS and GSM8K), where the attention baseline performs comparably to random noise, indicating its failure to capture essential reasoning structures.
+In contrast, GraphGhost maintains high fidelity even under low sparsity constraints.
+Moreover, on the challenging ARC dataset, GraphGhost demonstrates superior robustness; notably, it is the only method that surpasses the 0.5 fidelity threshold on Qwen3.
+These results suggest that while attention primarily captures surface-level correlations, GraphGhost is able to trace deeper, multi-hop structural dependencies that are critical for complex reasoning.
+4.3 Behaviors Controlled by Critical Nodes in GraphGhost: Dataset View
+To evaluate whether GraphGhost successfully identifies the most influential tokens in LLMs, we mute specific neuron nodes and observe the resulting changes in the outputs.
+Specifically, neuron nodes are muted by setting the corresponding positions of their layer–token pairs to zero during the reasoning process.
+We evaluate the impact through both case studies and statistical analysis.
+Figure 12: The impact of tokens on various layers. We mute the period token ’.’ for each layer.
 Top1-10 Top11-20
@@
 Table 1: The impacts of top tokens. “MAX” denotes the maximum influential token among the top-K slots
-We further analyze the relationship between the in-degree ratio across layers and the final reasoning predictions in the Qwen3-0.6B model, as illustrated in Figure 12.
+We first analyze the relationship between the in-degree ratio across layers and the final reasoning predictions in the Qwen3-0.6B model, as illustrated in Figure 12.
 Neuron nodes with higher in-degree values are likely to merge semantic information from multiple sources.
-To verify this, we compute the language-changing ratio when muting individual neuron nodes across layers, defined as $R=\frac{\#\mathrm{non-English\ answer}}{\#\mathrm{All\ answer}}$.where a non-English answer refers to any generated response that switches from English to another language after perturbation.
-As shown in Figure 12(a), the peaks of the language-changing ratio align closely with layers exhibiting high in-degree ratios.
-These layers are primarily distributed in the upper regions of the model, which also correspond to the increase in InFidelity scores observed in Figure 12(b).
-This pattern suggests that neuron nodes in higher layers are more likely to connect directly to logit neurons, thereby altering the logical flow of reasoning (e.g., causing generation loops) rather than simply substituting equivalent semantic representations.
-Figure 12: The impact of tokens on various layers. We mute the period token ’.’ for each layer.
+To verify this, we compute the language-changing ratio when muting individual neuron nodes across layers, defined as $R=\frac{\#\mathrm{non-English\ answer}}{\#\mathrm{All\ answer}}$, where a non-English answer refers to any generated response that switches from English to another language after perturbation.
+As shown in Figure 12(a), the peaks of the language-changing ratio align closely with layers exhibiting high in-degree ratios in middle layers.
+Furthermore, as shown in Figure 12(b), Infidelity increases in the upper layers.
+This indicates that neuron nodes in higher layers are more likely to connect directly to logit neurons, thereby altering the underlying reasoning logic.
+This observation is consistent with the findings in Section 3.2.2, where perturbations in middle layers primarily affect semantic expression, whereas perturbations in higher layers tend to modify the reasoning logic itself.
+We also evaluate the important tokens ranked by PageRank algorithm.
+We calculate the average InFidelity score when muting the corresponding neuron nodes and identify the maximum-impact neuron nodes within the top-$k$ positions.
+The results, summarized in Table 1, include the top-1–10 and top-11–20 nodes.
+Overall, the top-10 nodes show much higher influence, indicating that PageRank effectively identifies important nodes in GraphGhost.
+Larger models are more stable than smaller ones; Llama-8B achieves an InFidelity score of 0.28, whereas Qwen3 reaches 0.58.
+Notably, the <space> token is the most influential, especially in higher layers.
 5 Conclusion
-In this work, we introduced GraphGhost, a framework that models neuron activations and their propagation as graphs, providing a structural view of LLM reasoning.
-Using graph algorithms and structural interventions, we show how key neurons shape the logical flow and semantic consistency.
-These findings establish GraphGhost as an interpretation framework for understanding the structural foundations of reasoning in LLMs.
+We present GraphGhost, a graph-based framework that reveals the internal structures underlying LLM reasoning.
+By linking token interactions and neuron activations to graph properties, GraphGhost shows how semantic merging and reasoning triggers shape model behavior.
+Our results demonstrate that LLM reasoning is governed by structured internal graphs, offering a principled path toward mechanistic interpretability.
 Limitations
@@
 Advancing the design of circuit tracer would further enhance GraphGhost’s ability to identify key neuron nodes and capture more precise reasoning structures.
-Appendix A Related work
-A.1 Interpretation
-Interpreting how LLMs perform various reasoning tasks is an important step toward aligning them with human understanding Yao et al.
-(2023b); Zhao et al.
-(2024).
-Current work mainly falls into two directions: self-explanation Gu et al.
-(2024); Koa et al.
-(2024) and uncovering the underlying causal mechanisms Bloom et al.
-(2024); Dunefsky et al.
-(2024).
-The self-explanation approach, such as Chain-of-Thought prompting Wei et al.
-(2022); Kojima et al.
-(2022), requires LLMs to verbalize their reasoning steps, which can help guide them to correct answers.
-However, there is ongoing debate about whether such reasoning chains are truly faithful Turpin et al.
-(2023); Barez et al.
-(2025).
-A.2 Reasoning tasks
-Reasoning tasks are among the most important benchmarks for evaluating the ability of LLMs to solve complex problems.
-Specifically, they assess model performance across mathematical, scientific, and logical reasoning.
-Mathematical reasoning evaluates a model’s capacity to perform structured, multi-step numerical computations expressed in natural language Cobbe et al.
-(2021); Hendrycks et al.
-(2021).
-Scientific reasoning tasks require LLMs to draw on background knowledge and apply it in reasoning processes Khot et al.
-(2020); Clark et al.
-(2018).
-Logical reasoning tasks assess whether models can apply formal rules rather than relying on surface-level correlations Clark et al.
-(2019); Saparov and He (2022).
-Recent research further suggests that LLMs are capable of transferring reasoning strategies across diverse tasks, indicating that despite differences in domain, the underlying reasoning logic may be shared.
-Appendix B Algorithm
-Here we present the algorithm of construct the GraphGhost in Algorithm 1.
-Algorithm 1 GraphGhost Construction
-1:Input: A set of samples $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$, where each $s_{i}$ contains an answer token $t_{n}^{(i)}$
-2:Initialize global GraphGhost $G=(V,E,W)$, $V=\emptyset$, $E=\emptyset$, $W=\emptyset$
-3:for each sample $s_{i}\in\mathcal{S}$ do
-4:  Initialize token list $T=\{t_{n}^{(i)}\}$
+Potential Risk
+A potential risk of our method is that the extracted graph structures may be over-interpreted as definitive explanations of model behavior, while they only provide approximate views of internal information flow.
+We caution against such misuse and position our approach as a complementary analytical tool.
+Appendix A Appendix: Algorithm
+Here we present the algorithm of construct the sample and dataset view of GraphGhost in Algorithm 1 and Algorithm 2.
+Algorithm 1 GraphGhost Construction (Sample View)
+1:Input: A sample $s$ containing a final answer token $t_{n}$
+2:Output:
+  • Sample-level GraphGhost $G^{\text{sample}}=(V,E)$ (unweighted)
+  • A local edge-weight map $W^{\text{local}}$ (for dataset aggregation only)
+3:Initialize $V=\emptyset$, $E=\emptyset$, and $W^{\text{local}}(e)=0$ for all edges $e$
+4:Initialize frontier $T=\{t_{n}\}$
 5:for each token $t\in T$ do
-6:   if Token $t$ not in question of $s_{i}$ then
-7:     # trace contributing tokens and local attribution graph
-8:     $\hat{T},G_{A}^{\prime}\leftarrow\text{CircuitTracer}(t)$
+6:  if $t$ is not part of the question tokens then
+7:   # Step 1: Local logit-based attribution graph
+8:   $(\hat{T},G_{A}^{\prime})\leftarrow\textsc{CircuitTracer}(t)$ $\triangleright$ $G_{A}^{\prime}=(\hat{V},\hat{E},W^{\text{loc}})$
 9:   $T\leftarrow T\cup\hat{T}$
-10:     # update global GraphGhost with local graph
-11:     $\hat{V}=\{v^{1}(t_{1}),\dots,v^{\ell}(t_{n})\},\hat{E}\leftarrow G_{A}^{\prime}$
-12:     $V\leftarrow V\cup\hat{V}$
-13:     for each edge $e$ in $\hat{E}$ do
-14:      if $e\notin E$ then
-15:        $E\leftarrow E\cup\{e\}$
-16:        $W(e)\leftarrow 1$
-17:      else
-18:        $W(e)\leftarrow W(e)+1$
-19:      end if
-20:     end for
-21:   end if
-22:  end for
-23:end for
-24:return $G=(V,E,W)$
-Appendix C Appendix: Experiment setting
-C.1 Synthetic Experiment
+10:   # Step 2: Add edges to sample-level graph (structure only)
+11:   $V\leftarrow V\cup\hat{V}$
+12:   $E\leftarrow E\cup\hat{E} \triangleright$ unweighted sample graph
+13:   # Step 3: Accumulate local weights for dataset view
+14:   for each edge $e\in\hat{E}$ do
+15:     $W^{\text{local}}(e)\leftarrow W^{\text{local}}(e)+W^{\text{loc}}(e)$
+16:   end for
+17:  end if
+18:end for
+19:return $(G^{\text{sample}}=(V,E),\;W^{\text{local}})$
+Algorithm 2 GraphGhost Construction (Dataset View)
+1:Input: A set of samples $\mathcal{S}=\{s_{1},\dots,s_{m}\}$
+2:Output: Dataset-level GraphGhost $G^{\text{data}}=(V^{\text{data}},E^{\text{data}},W^{\text{data}})$
+3:Initialize $V^{\text{data}}=\emptyset$, $E^{\text{data}}=\emptyset$, and $W^{\text{data}}(e)=0$
+4:for each sample $s_{i}\in\mathcal{S}$ do
+5:  $(G^{(i)},W^{\text{local}})\leftarrow\textsc{GraphGhost-SampleView}(s_{i})$
+6:  $(V^{(i)},E^{(i)})\leftarrow G^{(i)}$
+7:  # Merge sample graph structure
+8:  $V^{\text{data}}\leftarrow V^{\text{data}}\cup V^{(i)}$
+9:  $E^{\text{data}}\leftarrow E^{\text{data}}\cup E^{(i)}$
+10:  # Accumulate edge weights across samples
+11:  for each edge $e\in E^{(i)}$ do
+12:   $W^{\text{data}}(e)\leftarrow W^{\text{data}}(e)+W^{\text{local}}(e)$
+13:  end for
+14:end for
+15:return $G^{\text{data}}=(V^{\text{data}},E^{\text{data}},W^{\text{data}})$
+Appendix B Appendix: Experiment setting
+B.1 Synthetic Experiment
 We first generate directed graphs with 10 nodes and an edge density of 0.6, then randomly drop 60% of the edges.
@@
 For training, we use a 5-layer decoder-only transformer with a hidden size of 96 and 12 attention heads.
-The model is trained for 20,000 iterations to ensure stable convergence.
+The model is trained for 20,000 iterations to ensure stable convergence where the validation loss is the minimal.
 For circuit tracing, we employ the ReLU version of the transcoder, with the dead feature window set to 50 and the L1 coefficient set to 0.0005.
 During visualization, we set the node ratio to 0.48 and the edge ratio to 0.86.
-C.2 Real-world Experiment
+B.2 Real-world Experiment
 The experimental settings for real-world data are summarized in Table 2.
@@
 Table 2: The experiment setting for real world cases. “Let’s…” denotes the prompt “Let’s think step-by-step”, and “<>” denotes “<think>”
-Appendix D GraphGhost Results
+B.3 Case studies
+We show the full case studies of how neuron nodes change the LLMs behaviors in Figure 14 and Figure 13.
+We show that muting neurons in middle layers leads to semantic changes, as illustrated in Figure 13.
+When the neuron node associated with “the” is deleted at layer 20, the English semantic pathway is disrupted, activating Chinese neurons that share similar semantic signals.
+As a result, the model produces output in Chinese, effectively translating from the original English text.
+Furthermore, in the word-expression change case, deleting the neuron associated with “.” at layer 17 causes the token “/” to replace “÷”.
+In Figure 14, we demonstrate how GraphGhost influences reasoning by altering key steps or driving the generation into loops.
+In the loop case, deleting the neuron node “the” at layer 27 causes the reasoning to degenerate into a repetition, producing “theShaper Information.” These cases suggest that even a single neuron in the middle layer can fundamentally change the outcome of the reasoning process.
+In the key-step change case, deleting the neuron node “.” at layer 25 introduces a new reasoning token such as “So", ultimately leading the model to produce the correct answer.
+Change the semantic Change the language Question: A ) nervous and digestive B ) nervous and circulatory C ) respiratory and digestive D ) respiratory and circulatory Question: Which two body systems would most directly remove extra fluid from a person’s lungs?
+Original answer: … B) Nervous and circulatory systems: The nervous system is responsible for regulating the body’s response to stress, while the circulatory system is responsible for transporting blood throughout the body.
+Neither of these systems directly remove extra fluid from the lungs… Mute the neuron node ’the’ at layer 20 Perturbed answer: … B) Nervous and circulatory systems: The nervous system is responsible for controlling the - circulatory system is responsible for transporting blood throughout the分区肺.
+While the分区肺需要血液来清除多余的水分,这些系统本身并不直接参与这个过程。…   Change the word expression Question: To determine how many trips Jerry will make, we need to calculate the total number of trays he needs to pick up and then divide by the number of trays he can carry at a time.
+Original answer: … we divide the total number of trays, 16.0 trays ÷ 8.0 trays per trip = 2 trips… Mute the neuron node ’.’ at layer 17 Perturbed answer: … Number of trips = Total trays / 8 = 16 / 8 = 2 trips …
+Figure 13: Semantic control in middle layers modulates changes in semantic representations. In the language-switching case, the generated characters are direct translations of the original answer.
+Change the reasoning structure Failed into loop Question: Alyssa picked 25.0 limes and Mike picked 32.0 limes.
+Tom picked 12.0 plums.
+How many limes were picked in all?
+Let’s think step by step Original answer: Step 1: Identify the number of limes picked by Alyssa and Mike.
+Alyssa picked 25.0 limes.
+Mike picked 32.0 limes.
+… Mute the neuron node ’the’ at layer 27 Perturbed answer: Step 1: Identify theSharper Information In this problem, we are given theSharper Information about theSharper Information about Alyssa, Mike, and Tom’s limes and plums.
+We need to find theSharper Information about theSharper Information about theSharper …   Change the key steps Question: There are 7.0 crayons in the drawer and 6.0 crayons on the desk.
+Sam placed 4.0 crayons and 8.0 scissors on the desk.
+How many crayons are now there in total?
+Let’s think step by step (ground truth: 17) Original answer: … - Sam placed 4.0 crayons on the desk.
+- Sam also placed 8.0 scissors on the desk.
+….
+- Total crayons on the desk = 7.0 + 4.0 + 8.0 - Total crayons on the desk = 19.0 Mute the neuron node ’.’ at layer 25 Perturbed answer: … - Sam placed 4.0 crayons on the desk.
+- So, the total number of crayons on the desk is now 6.0 (initial crayons) + 4.0 (Sam’s crayons) = 10.0 crayons.
+… - So, the total number of crayons = 7.0 (drawer) + 10.0 (desk) = 17.0 crayons.
+Figure 14: Reasoning triggers changes on the reasoning structures
+B.4 Attention-based Baseline Implementation
+To adapt the standard attention analysis (Clark et al., 2019) for reasoning tasks, we implement a specific Answer-to-Input saliency measure.
+We first calculate the mean attention matrix $\bar{\mathbf{A}}$ by averaging weights across all heads in the final layer.
+The importance score $S_{i}$ for each input token $i$ is then derived by accumulating the attention it receives from the entire generated answer sequence:
+$S_{i}=\sum_{j\in\text{Answer}}\bar{\mathbf{A}}_{j,i}$
+This score represents the total attention mass focused on input $i$ during the reasoning generation.
+Input tokens are ranked by $S_{i}$, and the top-$k\%$ are retained to construct the perturbed prompt for fidelity evaluation, following the same protocol as GraphGhost.
+Appendix C GraphGhost Results
 In this section, we present the specific tokens ranked by graph algorithms.
@@
 Table 7: The examples of how LLMs merge the semantics.
+Appendix D Additional Quantitative Results
+Figure 15 presents the detailed fidelity-sparsity curves for all evaluated models (DS-Qwen-1.5B, Qwen3-0.6B, and Qwen2.5-1.5B) across the three datasets.
+Unlike the aggregated results in the main text (Figure 11), this figure illustrates the individual performance trajectories for each model without averaging.
+Figure 15: Detailed fidelity-sparsity curves for individual models across ARC, MAWPS, and GSM8K datasets.
+This figure shows the non-averaged performance for GraphGhost compared to Random and Attention baselines.